### v1

# Geometric Calibration and Neutral Zones for Uncertainty-Aware Multi-Class Classification

Abstract
Modern artificial intelligence systems make critical decisions yet often fail silently when uncertain. We develop a geometric framework for post-hoc calibration of neural network probability outputs, treating probability vectors as points on the (c1)(c-1)-dimensional probability simplex equipped with the Fisher–Rao metric. Our approach yields Additive Log-Ratio (ALR) calibration maps that reduce exactly to Platt scaling for binary problems (Proposition 1) while extending naturally to multi-class settings—providing a principled generalization that existing methods lack. Complementing calibration, we define geometric reliability scores based on Fisher–Rao distance and construct neutral zones for principled deferral of uncertain predictions.
Theoretical contributions include: (i) consistency of the calibration estimator at rate $O_{p}(n^{-1/2})$ via M-estimation theory (Theorem 1), and (ii) tight concentration bounds for reliability scores with explicit sub-Gaussian parameters enabling sample size calculations for validation set design (Theorem 2). We conjecture Neyman–Pearson optimality of our neutral zone construction based on connections to Bhattacharyya coefficients. Empirical validation on Adeno-Associated Virus classification demonstrates that the two-stage framework (calibration followed by reliability-based deferral) captures 72.5% of errors while deferring 34.5% of samples. Notably, this operational gain is achievable with any well-calibrated probability output; the contribution of geometric calibration lies in its theoretical foundations rather than empirical superiority over simpler alternatives. This work bridges information geometry and statistical learning, offering formal guarantees relevant to applications requiring rigorous validation.
Keywords: Calibration, uncertainty quantification, Fisher-Rao geometry, manifold geometry, probability simplex, geometric reliability scores, neural networks, neutral zones
Preprint Note: This manuscript is under preparation for journal submission. Section 5 presents open problems where we invite community collaboration. Comments welcome at soumojit.das@wsu.edu.
1 Introduction
Modern AI systems make critical decisions in medicine, criminal justice, and autonomous systems, yet they provide no reliable mechanism to identify when predictions are likely wrong. A medical AI diagnosing cancer from imaging may output $\mathbf{p}=(0.51,0.49)$ for malignant vs. benign with the same apparent confidence as $\mathbf{p}=(0.99,0.01)$, despite the former representing profound uncertainty. This forced classification without uncertainty quantification leads to silent failures—the system appears to function normally while making unreliable predictions with potentially severe consequences.
The fundamental issue is miscalibration: neural network probability outputs do not reflect true likelihood of correctness. Guo et al. [1] demonstrated that modern deep neural networks exhibit Expected Calibration Error (ECE) often exceeding 10% on standard benchmarks. This manifests as systematic overconfidence, where networks assign high probabilities to misclassified samples [2], and complete absence of uncertainty awareness [3].
1.1 Motivation: High-Stakes Applications
The deployment of AI in high-stakes domains reveals a critical gap: while these systems achieve impressive accuracy, they fail catastrophically when uncertain. Three domains illustrate the urgency:
Medical imaging: CNN models classify CT scans for lung cancer detection [4, 5] and FMRI scans for brain disorders. These methods lack precision measures for individual predictions, yet clinical decisions require knowing when to consult human experts.
Gene therapy: Adeno-Associated Viruses (AAVs) are critical vectors where misclassification of cargo types (Empty, ssDNA, dsDNA) can trigger immune responses in patients. Our preliminary work shows 16.45% error rates without uncertainty quantification.
Social decisions: AI increasingly informs decisions about individuals in employment, education, and criminal justice [6, 7]. When humans are the decision-making unit, misclassification costs are socially significant and often irreversible.
In each domain, the cost of misclassification vastly exceeds the cost of deferring uncertain cases to human review. Current methods provide no principled mechanism for this deferral.
1.2 Existing Approaches and Limitations
Existing post-hoc calibration methods—temperature scaling [1] and Platt scaling [8]—apply univariate transformations that treat probability vectors as unconstrained Euclidean points, ignoring the simplex constraint $\sum_{i}p_{i}=1$. These methods achieve good calibration on average but provide no instance-level uncertainty quantification.
Uncertainty quantification approaches such as conformal prediction [9] provide coverage guarantees but produce set-valued predictions that are difficult to interpret. Bayesian neural networks [10] and deep ensembles [11] require significant computational overhead, making them impractical for many real-time applications.
Neutral zone methods have shown promise in binary classification. The r-power framework [12] identifies ambiguous predictions through hypothesis testing but extension to multi-class problems requires new theoretical machinery, as the binary hypothesis testing framework does not generalize directly.
1.3 Our Approach and Contributions
We treat neural network probability outputs as points on the (c1)(c-1)-dimensional probability simplex
$\Delta^{c-1}=\left\{\mathbf{p}\in\mathbb{R}^{c}:p_{i}\geq 0,\sum_{i=1}^{c}p_{i}=1\right\},$ (1)
a Riemannian manifold equipped with the Fisher-Rao metric (Definition 1). This geometric perspective reveals that existing calibration methods—which apply transformations treating probability vectors as unconstrained Euclidean points—fundamentally misspecify the problem by ignoring the curved structure of probability space.
Central theoretical result.
Our main contribution is establishing that calibration via Additive Log-Ratio (ALR) transformations reduces exactly to Platt scaling for binary problems (Proposition 1). This positions geometric calibration as the natural multi-class generalization of a widely-used method, providing principled justification where none previously existed. The ALR framework respects the compositional structure of probability vectors [13] while enabling standard optimization techniques in the transformed space.
Statistical guarantees.
We establish two theoretical results supporting the framework:
  1. 1.
    Consistency (Theorem 1): The calibration estimator converges to the population optimum at rate $O_{p}(n^{-1/2})$, with strong convexity ensured by regularization and additional curvature from the softmax Hessian structure.
  2. 2.
    Concentration (Theorem 2): Reliability scores based on Fisher-Rao distance satisfy tight concentration bounds with explicit sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$. These bounds enable sample size calculations for validation set design prior to data collection.
Geometric neutral zones.
Complementing calibration, we define reliability scores based on Fisher-Rao distance to the predicted simplex vertex (Definition 3) and construct neutral zones for principled deferral of uncertain predictions. The connection to Bhattacharyya coefficients (Remark 6) suggests potential Neyman-Pearson optimality, which we state as Conjecture 1.
Importantly, the neutral zone mechanism is method-agnostic: it can be applied to any calibration method. Our empirical analysis (Section 4.8, Table 5) confirms that operational gains are similar across calibration approaches. The contribution of geometric calibration is providing theoretical foundations, not empirical superiority.
Empirical validation.
On Adeno-Associated Virus (AAV) classification for gene therapy ($c=3$ classes, $n_{\text{val}}=310$), the two-stage framework achieves:
  • 72.5% of errors captured while deferring 34.5% of samples
  • Automated decision error rate reduced from 16.8% to 6.9%
  • Validation sample size ($n=310$) providing $\approx 5\times$ safety margin over theoretical requirement from Theorem 2
The modest scale of this application is appropriate for proof of concept; validation on larger-scale and more diverse settings remains future work (Section 5).
1.4 Novelty
This work makes the following contributions:
Mathematical framework:
To our knowledge, this is the first systematic application of Fisher-Rao geometry to post-hoc neural network calibration. While information geometry has been studied in the context of natural gradient descent and exponential families [14], its use for correcting probability outputs after training is novel. The key result—that geometric calibration reduces to Platt scaling for $c=2$ (Proposition 1)—validates our approach as a principled generalization rather than an arbitrary alternative.
Theoretical guarantees:
Theorem 2 provides the first concentration bounds for geometric reliability scores with explicit sub-Gaussian parameters. Previous work on uncertainty quantification (conformal prediction [9], deep ensembles [11]) either lacks sample complexity bounds or provides only asymptotic guarantees. Our explicit constants (Remark 10) enable practitioners to compute required validation set sizes.
Unified two-stage framework:
We decompose uncertainty-aware classification into: (i) calibration (population-level probability correction, Section 2.2) and (ii) reliability scoring (instance-level uncertainty quantification, Section 4.5). This separation clarifies that calibration alone provides modest accuracy gains (Table 4), while the combination with neutral zones (Section 4.6) achieves substantial error reduction. Our empirical analysis (Section 4.8, Table 5) demonstrates this is true regardless of calibration method, isolating the neutral zone mechanism’s contribution.
Information-theoretic grounding:
The connection between Fisher-Rao reliability scores and Bhattacharyya coefficients (Remark 6) links our geometric construction to classical signal detection theory [15] and Chernoff bounds [16]. This bridges modern geometric methods with established decision theory, suggesting potential optimality properties (Conjecture 1).
1.5 Paper Organization
Section 2 develops the mathematical framework, including geometry of the probability simplex, geometric calibration, and geometric neutral zones. Section 3 presents proven theorems and conjectures with supporting evidence. Section 4 provides empirical validation on AAV data. Section 5 discusses open problems and future directions.
2 Mathematical Framework
2.1 Preliminaries: Geometry of the Probability Simplex
Let $\mathbf{p}^{\text{CNN}}\in\Delta^{c-1}$ denote the probability vector output by a neural network for a $c$-class problem. The simplex admits a natural Riemannian structure.
Definition 1 (Fisher-Rao Metric).
The Fisher information metric at $\mathbf{p}\in\Delta^{c-1}$ is
$g_{ij}(\mathbf{p})=\frac{\delta_{ij}}{4p_{i}}$ (2)
where $\delta_{ij}$ is the Kronecker delta. This induces the Fisher-Rao distance:
$d_{\text{FR}}(\mathbf{p},\mathbf{q})=2\arccos\left(\sum_{i=1}^{c}\sqrt{p_{i}q_{i}}\right).$ (3)
Remark 1 (Geometric Foundations).
The Fisher-Rao metric arises as the unique Riemannian metric on the simplex (up to scaling) that is invariant under both coordinate permutation and reparameterization [14]. This invariance reflects the fundamental nature of probability vectors as compositional data [13]: only ratios between probabilities carry information, not their absolute magnitudes. The simplex constraint $\sum_{i}p_{i}=1$ represents arbitrary scaling.
From information geometry, the Fisher-Rao metric is the natural Riemannian structure induced by the categorical distribution as an exponential family. The $1/p_{i}$ terms in the metric arise from the variance structure of the sufficient statistics. The simplex under this metric has constant positive sectional curvature $K=+1/4$ and bounded diameter $\operatorname{diam}(\Delta^{c-1})=\pi$. This follows from the isometry between $(\Delta^{c-1},g)$ and the positive orthant of the sphere of radius $2$ in $\mathbb{R}^{c}$ via the square-root embedding $\mathbf{p}\mapsto 2(\sqrt{p_{1}},\ldots,\sqrt{p_{c}})$; since a sphere of radius $r$ has sectional curvature $K=1/r^{2}$, we obtain $K=+1/4$ [17, 14]. We exploit the bounded diameter in Theorem 2. The positive curvature is geometrically favorable: geodesic balls on spheres have smaller volume than Euclidean balls of the same radius, implying that probability mass concentrates more tightly around means. This spherical concentration (Lévy-type) strengthens the Hoeffding bounds in Theorem 2.  [^1] We emphasize that this positive curvature refers to the probability simplex $\Delta^{c-1}$ itself, viewed as the sample space of categorical distributions. This differs from the parameter space of the Dirichlet family, which has negative curvature [18]. The distinction matters: our reliability scores operate on probability outputs (elements of $\Delta^{c-1}$), not on distribution parameters.
Definition 2 (Additive Log-Ratio Transform).
The ALR transform [13] maps $\Delta^{c-1}$ to $\mathbb{R}^{c-1}$:
$\operatorname{ALR}(\mathbf{p})=\left(\log\frac{p_{1}}{p_{c}},\ldots,\log\frac{p_{c-1}}{p_{c}}\right)^{\top}$ (4)
with inverse $\operatorname{ALR}^{-1}(\mathbf{z})=\text{softmax}(z_{1},\ldots,z_{c-1},0)$.
Remark 2 (Coordinate Choice).
The ALR transform provides a coordinate system on the simplex that respects its compositional structure by expressing probabilities as log-ratios relative to a reference class (here, class $c$). This choice makes the problem scale-invariant: multiplying all probabilities by a constant leaves log-ratios unchanged. Alternative coordinates exist (e.g., isometric log-ratio, centered log-ratio), but ALR offers computational simplicity while preserving the essential geometry.
2.2 Geometric Calibration
2.2.1 Problem Formulation
A classifier is perfectly calibrated if $\mathbb{P}(Y=j\mid f(X)=\mathbf{p})=p_{j}$ for all classes $j$ and probability vectors $\mathbf{p}$. Given labeled training data $\{(\mathbf{p}_{i}^{\text{CNN}},y_{i})\}_{i=1}^{n}$, we seek an affine transformation in ALR space:
$\mathbf{z}^{\text{cal}}=A\cdot\operatorname{ALR}(\mathbf{p}^{\text{CNN}})+\mathbf{b}$ (5)
where $A\in\mathbb{R}^{(c-1)\times(c-1)}$ is positive definite and $\mathbf{b}\in\mathbb{R}^{c-1}$. The calibrated probabilities are:
$\mathbf{p}^{\text{cal}}=\operatorname{ALR}^{-1}(A\cdot\operatorname{ALR}(\mathbf{p}^{\text{CNN}})+\mathbf{b}).$ (6)
Remark 3 (Model Specification).
Why restrict to affine transformations in ALR space? This choice balances statistical efficiency and model parsimony. In ALR coordinates, calibration becomes a regression problem: we model the relationship between uncalibrated log-ratios and true probabilities as linear with additive noise. The affine structure provides sufficient flexibility to capture common miscalibration patterns (overconfidence, class-specific bias) while remaining identifiable with limited calibration data.
More complex nonlinear transformations (e.g., neural networks, kernel methods) risk overfitting when calibration sets are small—typical in medical applications where labeled validation data is expensive. The linear model in ALR space corresponds to a specific nonlinear transformation on the simplex itself, offering a middle ground between univariate methods (temperature scaling) and fully nonparametric approaches.
From a measurement error perspective, we treat $\mathbf{p}^{\text{CNN}}$ as a noisy observation of the true probabilities, distorted by the training process. The transformation $(A,\mathbf{b})$ estimates the inverse of this distortion map. Our geometric approach posits that this inverse is approximately linear in the natural (ALR) coordinate system of the simplex.
2.2.2 Learning Objective
We minimize the regularized cross-entropy loss:
[…]
subject to $A\succeq\delta I$ for some $\delta>0$ and $\operatorname{tr}(A)=c-1$ for identifiability. Here $y_{ij}=\mathbf{1}[y_{i}=j]$ is the one-hot encoding.
From a statistical perspective, we model the uncalibrated CNN outputs as systematically biased observations of the true probabilities. The affine transformation $(A,\mathbf{b})$ corrects both multiplicative bias (through $A$) and additive bias (through $\mathbf{b}$) in ALR space. The regularization towards identity transformation reflects our prior belief that well-trained networks require minimal correction.
Remark 4 (Regularization and Prior Specification).
The regularization terms admit both frequentist and Bayesian interpretations. From an empirical risk minimization perspective: (i) $\|A-I\|_{F}^{2}$ prevents overfitting by penalizing unnecessary rotations and scaling in ALR space, (ii) $\|\mathbf{b}\|^{2}$ controls translation magnitude, and (iii) together they implement a “minimal intervention” principle—we seek the smallest correction that achieves calibration.
From a Bayesian perspective, our regularization corresponds to the prior $(A-I)\sim\mathcal{N}(0,\lambda_{1}^{-1}I)$ and $\mathbf{b}\sim\mathcal{N}(0,\lambda_{2}^{-1}I)$, centered on the identity map. This formalizes our belief that modern neural networks, despite miscalibration, are not arbitrarily wrong—the calibration map should be close to identity. The regularization strength $(\lambda_{1},\lambda_{2})$ quantifies this belief: larger values impose stronger shrinkage toward the null hypothesis “CNN is already calibrated.”
This prior specification is particularly important given limited calibration data. Medical imaging applications may have only $10^{2}$$10^{3}$ labeled validation samples, while the unconstrained parameter space has dimension $(c-1)^{2}+(c-1)$. The regularization effectively reduces the model complexity to match the available sample size.
2.3 Geometric Neutral Zones
Even after calibration, some predictions remain inherently uncertain. We identify these via geometric reliability scores.
Definition 3 (Reliability Score).
For calibrated probability $\mathbf{p}^{\text{cal}}$ with predicted class $\hat{j}=\operatorname*{arg\,max}_{j}p_{j}^{\text{cal}}$, the reliability score is:
[…]
where $\mathbf{e}_{j}$ is the $j$-th vertex of the simplex and $\lambda>0$ is a sensitivity parameter.
Remark 5 (Decision-Theoretic Interpretation).
The reliability score $R$ serves as a candidate sufficient statistic for the binary decision problem $\{\text{defer to human},\text{classify automatically}\}$. Under the monotone likelihood ratio property (Conjecture 1), no other function of $\mathbf{p}^{\text{cal}}$ improves the decision rule—all information relevant to the deferral decision is captured by the Fisher-Rao distance to the predicted vertex.
This dimension reduction from $\Delta^{c-1}$ (the full probability vector) to $\mathbb{R}_{+}$ (the scalar reliability score) is justified geometrically: for decision-making purposes, what matters is not the detailed structure of the probability vector, but rather how far it lies from certainty (the simplex vertices). The Fisher-Rao metric provides the “correct” notion of distance for this purpose, respecting the simplex geometry.
The exponential transformation $\exp(-\lambda\cdot d)$ converts distances to a $[0,1]$ reliability scale, with $\lambda$ controlling sensitivity. Larger $\lambda$ makes the score more discriminative between near-vertex (reliable) and near-center (unreliable) predictions.
Definition 4 (Neutral Zone).
The neutral zone at level $\alpha$ is:
$NZ_{\alpha}=\{i:R(\mathbf{p}_{i}^{\text{cal}})<\tau^{*}\}$ (9)
where $\tau^{*}$ satisfies $\mathbb{P}(\text{error}\mid R\geq\tau^{*})\leq\alpha$.
Remark 6 (Information-Theoretic Grounding).
[…]
This connection, extensively studied in signal detection [15], suggests our geometric neutral zones inherit optimality properties from classical decision theory. Specifically, if the reliability score induces a monotone likelihood ratio (Conjecture 1), then threshold tests on $R$ are Neyman-Pearson optimal: among all tests that defer at most $\alpha$ fraction of samples, our neutral zone maximizes the proportion of errors captured. This provides information-theoretic grounding for our geometric construction.
3 Theoretical Results
We present three categories of results: proven theorems (Section 3.1), conjectures with strong evidence (Section 3.2), and research directions (Section 5).
3.1 Proven Results
3.1.1 Reduction to Platt Scaling
Proposition 1 (Binary Case Reduction).
For $c=2$, geometric calibration reduces to Platt scaling. Specifically, if $\operatorname{ALR}(p)=\log\frac{p}{1-p}=\operatorname{logit}(p)$, then the calibration map becomes:
$p^{\text{cal}}=\sigma(a\cdot\operatorname{logit}(p^{\text{CNN}})+b)$ (10)
where $\sigma$ is the logistic function and $(a,b)\in\mathbb{R}^{2}$ are learned parameters.
Proof.
For $c=2$, let $p=p_{1}$ so that $p_{2}=1-p$. The ALR transform with reference class $c=2$ gives:
$\operatorname{ALR}(p)=\log\frac{p}{1-p}=\operatorname{logit}(p).$ (11)
The calibration map (5) becomes scalar: $z^{\text{cal}}=a\cdot\operatorname{logit}(p^{\text{CNN}})+b$ where $a\in\mathbb{R}_{>0}$ and $b\in\mathbb{R}$. The inverse ALR is:
$p^{\text{cal}}=\operatorname{ALR}^{-1}(z^{\text{cal}})=\frac{\exp(z^{\text{cal}})}{1+\exp(z^{\text{cal}})}=\sigma(z^{\text{cal}})=\sigma(a\cdot\operatorname{logit}(p^{\text{CNN}})+b).$ (12)
This is precisely the Platt scaling formulation [8]. ∎
Remark 7.
Proposition 1 validates our framework: it recovers the established binary method while providing a principled multi-class extension.
3.1.2 Consistency and Convergence Rate
Theorem 1 (Consistency of Geometric Calibration).
Let $\{(\mathbf{p}_{i}^{\text{CNN}},y_{i})\}_{i=1}^{n}$ be i.i.d. samples with $\mathbf{p}_{i}^{\text{CNN}}$ bounded away from the simplex boundary: $p_{ij}\geq\epsilon>0$ for all $i,j$. Let $\hat{T}_{n}=(\hat{A}_{n},\hat{\mathbf{b}}_{n})$ minimize (7) and let $T^{*}=(A^{*},\mathbf{b}^{*})$ minimize the population risk $\mathcal{L}^{*}(T)=\operatorname{\mathbb{E}}[\ell(T;\mathbf{p},y)]$. Then:
  1. (i)
    $\hat{T}_{n}\xrightarrow{p}T^{*}$ as $n\to\infty$.
  2. (ii)
    $\|\hat{T}_{n}-T^{*}\|=O_{p}(n^{-1/2})$.
Proof.
See Appendix A.1 for the complete proof.
Proof sketch:
  1. 1.
    Strong convexity: The Hessian of $\mathcal{L}$ satisfies $\nabla^{2}\mathcal{L}(T)\succeq\mu I$ where:
    $\mu=\min\{2\lambda_{1},2\lambda_{2}\}>0$ (13)
    The regularization terms provide uniform strong convexity. The data term contributes additional (but not uniform) curvature via the softmax Hessian structure (Lemma 2).
  2. 2.
    Compact parameter space: Regularization confines parameters to a compact set $\mathcal{T}$.
  3. 3.
    Uniform convergence: Bounded loss (proven via chain of bounds on ALR inputs, calibrated logits, and softmax outputs) yields Glivenko-Cantelli.
  4. 4.
    Apply M-estimation theory: By van der Vaart [20], Theorems 5.7 and 5.23, consistency and the $O_{p}(n^{-1/2})$ rate follow.
Remark 8 (Role of Regularization).
Strong convexity of $\mathcal{L}^{*}$ derives primarily from the regularization terms, with $\mu=\min\{2\lambda_{1},2\lambda_{2}\}$. The data term $\mathbb{E}[J_{T}^{\top}\Sigma_{\mathbf{p}^{\text{cal}}}J_{T}]$ is positive semi-definite and inherits curvature from the softmax Hessian (Lemma 2), but because the Jacobian $J_{T}$ maps to a (c1)(c-1)-dimensional space while the parameter space has dimension (c1)c(c-1)c, the data term alone cannot provide uniform strong convexity. This underscores the importance of regularization in finite-sample settings—it ensures identifiability and controls the convergence rate even when calibration data is limited.
3.1.3 Concentration of Reliability Scores
Theorem 2 (Reliability Concentration).
Let $)R(\mathbf{p})=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{e}_{\hat{j}}$)) as in Definition 3, where $\hat{j}=\operatorname*{arg\,max}_{k}p_{k}$. For $\mathbf{p}$ drawn from any distribution on $\Delta^{c-1}$:
  1. (i)
    Tail bound:
    $\mathbb{P}(|R-\operatorname{\mathbb{E}}[R]|>t)\leq 2\exp\left(-\frac{2t^{2}}{(1-e^{-\lambda\pi})^{2}}\right).$ (14)
  2. (ii)
    Sub-Gaussian parameter: $R-\operatorname{\mathbb{E}}[R]$ is sub-Gaussian with $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$.
  3. (iii)
    For $\lambda=1$: $\mathbb{P}(|R-\operatorname{\mathbb{E}}[R]|>t)\leq 2\exp(-2.18\,t^{2})$.
Proof.
See Appendix A.2 for the complete proof.
Proof sketch:
  1. 1.
    Continuity of $R$: The reliability score is continuous on all of $\Delta^{c-1}$. At decision boundaries where $p_{i}=p_{j}$, we have $d_{\text{FR}}(\mathbf{p},\mathbf{e}_{i})=2\arccos(\sqrt{p_{i}})=2\arccos(\sqrt{p_{j}})=d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j})$, so the $\operatorname*{arg\,max}$ switching introduces no discontinuity.
  2. 2.
    Bounded range: Since $\operatorname{diam}(\Delta^{c-1})=\pi$, we have $R\in[e^{-\lambda\pi},1]$.
  3. 3.
    Apply Hoeffding: The bounded range $(1-e^{-\lambda\pi})$ yields the concentration bound directly.
Remark 9.
[…]
Remark 10 (Practical Implications of Concentration Bounds).
The concentration bound in Theorem 2 has direct implications for validation set design. For a desired deviation probability $\delta$ and precision $t$, our bound requires $n\approx\sigma^{2}\log(2/\delta)/(2t^{2})$ samples to ensure $\mathbb{P}(|R-\mathbb{E}[R]|>t)\leq\delta$, where $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$ is the sub-Gaussian parameter. With $\lambda=1$, we have $\sigma^{2}\approx 0.23$, giving $n\approx 0.12\log(2/\delta)/t^{2}$.
For high-reliability applications requiring $\delta=0.01$ and precision $t=0.1$, this yields $n_{\text{ours}}\approx 61$ samples. A generic Lipschitz-diameter bound (treating the reliability function as having range $\lambda\pi$) would yield $\sigma^{2}_{\text{naive}}=(\lambda\pi)^{2}/4\approx 2.47$, requiring $n_{\text{naive}}\approx 654$ samples—an order-of-magnitude difference.
Two caveats apply: (i) practitioners who estimate the reliability score’s empirical range from pilot data would achieve similar efficiency to our bound, and (ii) the sample complexity for the full inferential procedure is higher since $\mathbb{E}[R]$ must also be estimated. The theoretical contribution is deriving these bounds a priori from the geometric structure, enabling sample size planning before data collection. Our empirical validation used $n=310$ samples, providing approximately $5\times$ safety margin above the theoretical minimum.
3.2 Conjectures
The following conjectures have strong empirical support but proofs remain in progress. We present them with explicit evidence and testable mechanisms.
3.2.1 Empirical Convergence Behavior
Bootstrap analysis on AAV data suggests the estimator may converge faster than the asymptotic $O_{p}(n^{-1/2})$ rate in finite samples, though this observation requires careful interpretation.
Remark 11 (Finite-Sample Convergence).
Log-log regression of calibration error versus sample size yields slope 0.82-0.82 (95% CI: [1.11,0.52][-1.11,-0.52]); see Figure 1. Several caveats temper this observation:
  1. 1.
    Target mismatch: The analysis measures convergence $\hat{T}_{n}\to\hat{T}_{1241}$ (distance to the empirical optimum), not $\hat{T}_{n}\to T^{*}$ (distance to the population parameter). Convergence to a finite-sample estimate can appear faster than convergence to the true parameter.
  2. 2.
    Marginal significance: The confidence interval lower bound 0.52-0.52 is statistically indistinguishable from the asymptotic rate $\alpha=0.5$. The evidence for $\alpha>0.5$ is suggestive but not conclusive.
  3. 3.
    Regularization effects: The apparent acceleration may reflect reduced effective dimension due to strong regularization rather than intrinsic geometric properties.
We present this as an empirical observation warranting further investigation—particularly simulation studies with known population parameters—rather than a formal conjecture. The theoretical rate $O_{p}(n^{-1/2})$ established in Theorem 1 remains the rigorous guarantee.
Validation strategy for future work: To rigorously test whether super-efficiency occurs, we propose: (i) simulation studies with known $T^{*}$ where the true calibration map is specified, (ii) computing effective dimension via $d_{\text{eff}}=\operatorname{tr}(\nabla^{2}\mathcal{L}^{*})/\|\nabla^{2}\mathcal{L}^{*}\|_{2}$, and (iii) varying regularization strength $(\lambda_{1},\lambda_{2})$ to assess persistence of the phenomenon.
3.2.2 Neyman-Pearson Optimality
Conjecture 1 (NP Optimality of Geometric Neutral Zones).
The reliability score $)R=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{e}_{\hat{j}}$)) induces a monotone likelihood ratio ordering: for $R_{1}>R_{2}$,
$\frac{\mathbb{P}(\text{correct}\mid R_{1})}{\mathbb{P}(\text{error}\mid R_{1})}>\frac{\mathbb{P}(\text{correct}\mid R_{2})}{\mathbb{P}(\text{error}\mid R_{2})}.$ (15)
Consequently, threshold tests on $R$ are Neyman-Pearson optimal for the binary hypothesis test $H_{0}$: error vs. $H_{1}$: correct.
[…]
Path to proof:
  1. 1.
    Binary case ($c=2$): For two classes, the Fisher-Rao distance reduces to $d_{\text{FR}}(\mathbf{p},\mathbf{e}_{1})=2\arccos(\sqrt{p_{1}})$, which is a monotone function of $p_{1}$. Under a Dirichlet noise model $\mathbf{p}^{\text{CNN}}\sim\text{Dir}(\alpha\cdot\mathbf{p}^{\text{true}})$ with concentration $\alpha>0$, the conditional density $f(d_{\text{FR}}\mid Y=1)$ versus $f(d_{\text{FR}}\mid Y=0)$ can be computed explicitly via change of variables from the Beta distribution. The likelihood ratio $f(d_{\text{FR}}\mid Y=1)/f(d_{\text{FR}}\mid Y=0)$ is then a function of $d_{\text{FR}}$ alone. We are currently deriving this density to verify monotone likelihood ratio analytically for $c=2$.
  2. 2.
[…]
  3. 3.
    General proof: The main obstacle is characterizing the distribution of $d_{\text{FR}}$ under Dirichlet perturbations on $\Delta^{c-1}$. Two potential approaches: (a) derive explicit formulas for geodesic distances under wrapped normal or Riemannian normal distributions on the simplex–these are known for certain symmetric cases [14], or (b) use concentration of measure: the positive curvature $K=+1/4$ implies Lévy-type concentration on spherical caps, which may yield sufficient control on conditional densities to establish MLR via comparison arguments. Spherical concentration is actually stronger than hyperbolic concentration for our purposes, as geodesic balls on positively curved spaces have smaller volume than Euclidean balls of the same radius.
Validation strategy: Implement Monte Carlo validation for $c\in\{2,3,4,5\}$ with varying concentration parameters $\alpha\in\{1,5,10,50\}$. For each $(c,\alpha)$ pair, generate $10^{5}$ samples and estimate likelihood ratio $\widehat{LR}(d_{\text{FR}})=\hat{f}(d_{\text{FR}}\mid\text{correct})/\hat{f}(d_{\text{FR}}\mid\text{error})$ via kernel density estimation. Check monotonicity by testing whether $\partial\widehat{LR}/\partial d_{\text{FR}}<0$ holds across the support.
4 Empirical Validation
4.1 Experimental Setup
4.1.1 Application Domain
Adeno-Associated Viruses (AAVs) are critical vectors for gene therapy, delivering therapeutic genetic material to target cells. Misclassification of AAV cargo types—Empty (no genetic payload), single-stranded DNA (ssDNA), or double-stranded DNA (dsDNA)—can trigger immune responses in patients or result in manufacturing losses exceeding \$500,000 when contaminated batches require destruction. Current quality control relies on labor-intensive electron microscopy, creating a bottleneck in therapeutic development.
We applied our geometric calibration framework to a convolutional neural network trained on electrical signals from solid-state nanopore measurements [21, 22]. Previous work using r-power neutral zones on binary AAV classification achieved approximately 50% error capture while deferring 20% of samples [12]. Our geometric approach extends this methodology to the full 3-class problem with substantially improved performance.
4.1.2 Data and Model Architecture
Dataset: CNN trained on $n_{\text{train}}=1241$ nanopore signal traces, with performance evaluated on $n_{\text{val}}=310$ independent samples. Class distribution in the validation set: Empty ($n=131$, 42.3%), ssDNA ($n=57$, 18.4%), dsDNA ($n=122$, 39.4%). The validation set maintains similar proportions to training, ensuring representative evaluation across all cargo types.
Neural Network: ResNet-50 architecture pretrained on ImageNet, fine-tuned on AAV signal spectrograms. The network outputs unnormalized logits passed through softmax to produce probability vectors $\mathbf{p}^{\text{CNN}}\in\Delta^{2}$. Baseline accuracy: 83.2% on validation data, with 16.8% error rate concentrated in ssDNA classification (most difficult class due to signal ambiguity).
4.1.3 Calibration Procedure
Training phase:
  1. 1.
    Optimize loss function (7) via L-BFGS with regularization $\lambda_{1}=\lambda_{2}=0.01$
  2. 2.
    Validate positive definiteness: verify $\lambda_{\min}(A)>0$ throughout optimization
  3. 3.
    Set reliability parameter $\lambda=1.0$
  4. 4.
    Learn threshold $\tau^{*}=\sup\{\tau:\mathbb{P}(\text{error}\mid R\geq\tau)\leq\alpha\}$ on training set with $\alpha=0.05$
  5. 5.
    Selected hyperparameters: $\lambda=1.0$, $\tau^{*}=0.4451$
Learned calibration parameters:
$A=\begin{pmatrix}0.988&-0.002\\ -0.044&1.238\end{pmatrix},\quad\mathbf{b}=\begin{pmatrix}-0.017\\ 0.636\end{pmatrix}$ (16)
with $\operatorname{tr}(A)=2.23$ and $\lambda_{\min}(A)=0.987>0$, confirming positive definiteness.
Deployment protocol: For new sample with CNN output $\mathbf{p}^{\text{CNN}}$:
  1. 1.
    Calibrate: $\mathbf{p}^{\text{cal}}=\text{ALR}^{-1}(A\cdot\text{ALR}(\mathbf{p}^{\text{CNN}})+\mathbf{b})$
  2. 2.
    Compute reliability: $)R=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p}^{\text{cal}},\mathbf{e}_{\hat{j}}$)) where $\hat{j}=\operatorname*{arg\,max}_{k}p_{k}^{\text{cal}}$
  3. 3.
    Decision rule: If $R\geq\tau^{*}$, classify automatically; otherwise defer to human expert review
This procedure requires only forward passes through the calibration map and Fisher-Rao distance computation, adding negligible overhead (<<1ms per sample) to the CNN inference time.
4.2 Convergence Analysis
Before evaluating classification performance, we validate Remark 11: that geometric calibration converges faster than the theoretical $O_{p}(n^{-1/2})$ rate from Theorem 1.
4.2.1 Bootstrap Methodology
We assess finite-sample stability by measuring calibration discrepancy $\|\hat{T}_{n}-\hat{T}_{1241}\|$ as a function of training set size $n$, where $\hat{T}_{1241}=(\hat{A}_{1241},\hat{\mathbf{b}}_{1241})$ serves as a reference learned from the full training set. For each subsample size $n\in\{100,250,500,750,1000\}$:
  1. 1.
    Draw $n$ samples without replacement from the 1,241 training samples
  2. 2.
    Fit calibration parameters $\hat{T}_{n}=(\hat{A}_{n},\hat{\mathbf{b}}_{n})$ via (7) with identical hyperparameters
  3. 3.
    Compute Frobenius distance: $\text{Error}_{n}=\sqrt{\|\hat{A}_{n}-\hat{A}_{1241}\|_{F}^{2}+\|\hat{\mathbf{b}}_{n}-\hat{\mathbf{b}}_{1241}\|^{2}}$
  4. 4.
    Repeat 1,000 times to obtain distribution of $\text{Error}_{n}$
4.2.2 Results
Figure 1 presents convergence analysis on both linear and log-log scales.
Figure 1: Finite-sample convergence of geometric calibration. (Left) Calibration discrepancy $\|\hat{T}_{n}-\hat{T}_{1241}\|$ versus sample size $n\in\{100,250,500,750,1000\}$ with 1,000 bootstrap iterations per size. Blue points: empirical mean $\pm$ 1 standard deviation; red dashed line: theoretical $O(n^{-1/2})$ curve. Error decreases faster than theoretical prediction. (Right) Log-log plot reveals linear relationship with slope $\alpha=-0.82$ (95% CI: [1.11,0.52][-1.11,-0.52]), providing evidence that finite-sample convergence exceeds the asymptotic rate. The confidence interval excludes $\alpha=-0.5$, though the lower bound 0.52-0.52 is close to this threshold.
Observed convergence rate: Log-log regression yields slope $\alpha=-0.82$ with 95% confidence interval [1.11,0.52][-1.11,-0.52], providing evidence that finite-sample convergence exceeds the asymptotic $O_{p}(n^{-1/2})$ rate.
Variance reduction: Standard deviation of bootstrap estimates decreases from 0.20 at $n=100$ to 0.03 at $n=1000$, indicating stable parameter recovery even at moderate sample sizes.
Table 1: Bootstrap convergence summary ($B=1000$ iterations per sample size).
$n$ Mean Error SD
100 0.555 0.202
250 0.323 0.119
500 0.193 0.070
750 0.128 0.046
1000 0.077 0.028
4.2.3 Interpretation
The observed super-efficiency likely arises from structure in the parameter space that reduces the effective degrees of freedom. For $c=3$ classes, the nominal dimension is $(c-1)^{2}+(c-1)=6$, but several factors impose constraints: positive definiteness ($A\succeq\delta I$), trace normalization ($\operatorname{tr}(A)=c-1$), and the regularization prior centered on the identity transformation. Moreover, if the covariance structure of $\text{ALR}(\mathbf{p}^{\text{CNN}})$ concentrates on a low-dimensional subspace, the Hessian $\nabla^{2}\mathcal{L}^{*}(T^{*})$ may have only a few dominant eigenvalues, effectively reducing the estimation problem’s complexity. Characterizing the precise relationship between this effective dimension and the convergence rate remains an open problem.
Additionally, the softmax Hessian structure (Lemma 2) contributes favorable curvature in the directions most relevant to classification, while regularization ensures uniform strong convexity across the full parameter space. The interplay between the Fisher-Rao geometry (with sectional curvature $K=1/4$) and the ALR parameterization may contribute to the favorable loss landscape, though the precise mechanism requires further investigation.
4.3 Classification Performance
Figure 2 presents confusion matrix analysis comparing uncalibrated CNN outputs, geometrically calibrated outputs, and automated decisions outside the neutral zone.
Figure 2: Confusion matrices for AAV cargo classification ($n_{\text{val}}=310$). (Left) Uncalibrated CNN: 83.2% accuracy, 52 errors. (Center) After geometric calibration: 83.5% accuracy, 51 errors. (Right) Automated decisions outside neutral zone ($n=203$): 93.1% accuracy, 14 errors. Calibration alone yields marginal accuracy improvement; the substantial gain (6.9% error rate vs. 16.8% baseline) comes from uncertainty-aware deferral via neutral zones.
Key observation: Geometric calibration improves overall accuracy only marginally (83.2% $\to$ 83.5%, one additional correct prediction). The ssDNA class shows targeted improvement (71.9% $\to$ 77.2% recall), but this modest gain understates calibration’s value. The true benefit lies not in accuracy improvement, but in enabling reliable uncertainty quantification—the foundation for the neutral zone mechanism explored in Sections 4.54.6.
Table 2: Per-class performance metrics ($n_{\text{val}}=310$).
Class Precision Recall F1 NZ%
Empty .858 .878 .868 33.6
ssDNA .863 .772 .815 38.6
dsDNA .800 .820 .810 33.6
Macro Avg .840 .823 .831 34.5
Table 2 summarizes per-class performance. All classes achieve F1-scores $\geq 0.81$, with precision and recall within 10 percentage points—critical balance for medical applications. The ssDNA class enters the neutral zone at highest rate (38.6%), appropriately reflecting its lower baseline accuracy and greater inherent signal ambiguity.
4.4 Step 1: Probability Estimate Quality
Before assessing instance-level reliability (Step 2), we verify that geometric calibration improves the quality of probability estimates themselves.
4.4.1 Calibration Metrics
We evaluate calibration using three metrics: log loss and Brier score (strictly proper scoring rules) and Expected Calibration Error (ECE, a binned approximation of calibration quality). Table 3 presents per-class and overall results.
Table 3: Per-class and overall calibration metrics. Geometric calibration substantially improves the hardest class (ssDNA: $25\%$ log loss, $37\%$ ECE) while accepting modest degradation for already well-calibrated classes. Overall proper scoring rules improve; overall ECE increases (see Remark 12).
Class Metric Uncalibrated Calibrated Change
Empty Log Loss 0.311 0.319 $2.7\%$
Brier 0.178 0.183 $2.5\%$
ECE 0.023 0.030 $28\%$
ssDNA Log Loss 0.724 0.540 $\mathbf{-25.4\%}$
Brier 0.404 0.298 $\mathbf{-26.3\%}$
ECE 0.042 0.027 $\mathbf{-37\%}$
dsDNA Log Loss 0.407 0.448 $9.9\%$
Brier 0.236 0.257 $9.0\%$
ECE 0.045 0.037 $17\%$
Overall Log Loss 0.425 0.410 $\mathbf{-3.4\%}$
Brier 0.243 0.233 $\mathbf{-3.8\%}$
ECE 0.016 0.022 $33\%$
Interpretation. Geometric calibration performs targeted probability redistribution. The hardest class (ssDNA) improves dramatically across all metrics: log loss decreases by 25.4%, Brier score by 26.3%, and class-wise ECE by 37%. This reflects the ALR transformation learning to correct the CNN’s systematic under-prediction of ssDNA—the class with the highest error rate and lowest recall in the uncalibrated model.
The Empty class shows modest ECE degradation (+28%), reflecting the trade-off inherent in learning a single affine transformation for all classes. Importantly, the strictly proper scoring rules (Log Loss, Brier) improve overall, confirming that the probability estimates are genuinely better in an information-theoretic sense.
Remark 12 (Overall ECE Increase).
Two of three classes show per-class ECE improvement, yet overall ECE increases by 33%. This apparent paradox arises because the two metrics use different binning: per-class ECE bins samples by $p_{j}$ (predicted probability for class $j$), while overall ECE bins by $\max_{k}p_{k}$ (maximum probability across classes).
Geometric calibration performs selective sharpening: it increases confidence for correct predictions (mean max-prob: $0.872\to 0.885$, $p<0.001$ by paired $t$-test) while leaving confidence for errors essentially unchanged ($0.716\to 0.715$, $p=0.92$). This selective sharpening improves the reliability score’s discrimination ability (Section 4.5) at the cost of increased overall ECE. Specifically, the confidence bin $[0.5,0.6]$ contributes +0.0053+0.0053 to overall ECE (accounting for 99% of the increase), as samples move into this transition region where accuracy (46%) falls below confidence (55%).
For our application, the reliability score RR—not ECE—governs deferral decisions, making this trade-off favorable.
4.4.2 Reliability Diagrams
Figure 3 visualizes per-class calibration via reliability diagrams, comparing predicted probability against empirical frequency.
Figure 3: Per-class reliability diagrams ($n_{\text{val}}=310$, adaptive binning). Each panel plots mean predicted probability (x-axis) against empirical frequency (y-axis); perfect calibration follows the dashed diagonal. LOESS curves track the diagonal more closely after calibration, particularly for minority class ssDNA.
Remark 13 (Calibration Provides Population-Level Guarantees Only).
The metrics in Table 3 and Figure 3 confirm that Step 1 succeeds: geometric calibration produces better probability estimates. However, these are population-level guarantees. When we say “the model is well-calibrated,” we mean that among all predictions with 70% confidence, approximately 70% are correct on average.
This does not answer the operationally critical question: which specific predictions should we trust? A well-calibrated model may output $\mathbf{p}=(0.7,0.2,0.1)$ for both an “easy” sample and a “hard” sample. Calibration alone cannot distinguish these cases. Step 2 addresses this gap.
4.5 Step 2: Instance-Level Reliability Scoring
The reliability score $)R=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p}^{\text{cal}},\mathbf{e}_{\hat{j}}$)) translates calibrated probability vectors into scalar uncertainty measures. Unlike calibration metrics (which assess population-level properties), $R$ provides instance-level discrimination: high $R$ indicates predictions likely to be correct; low $R$ indicates predictions that warrant human review.
4.5.1 Score Distributions
Figure 4 presents reliability score distributions stratified by prediction outcome.
Figure 4: Reliability score distributions ($\lambda=1.0$). Correct predictions (blue): mean =0.579=0.579. Errors (red): mean =0.371=0.371. Separation $\Delta=0.208$ (effect size $d=1.17$). Dashed line: threshold $\tau^{*}=0.445$.
Correct predictions ($n=259$) concentrate at high reliability with mean $R=0.579$, while errors ($n=51$) concentrate at low reliability with mean $R=0.371$. The separation $\Delta=0.208$ corresponds to Cohen’s d=1.17d=1.17—a large effect size enabling effective discrimination. The distributions exhibit bimodal separation, consistent with Theorem 2: reliability scores concentrate around class-conditional means with sub-Gaussian parameter $\sigma^{2}\approx 0.23$.
4.5.2 Error Detection Performance
Figure 5 evaluates the reliability score as a binary classifier for error detection. We reduce the multi-class classification problem to a binary decision task: given a predicted class $\hat{j}$, is the prediction correct ($y=\hat{j}$) or incorrect ($y\neq\hat{j}$)? The reliability score $R$ serves as the decision criterion, with low $R$ indicating likely errors.
Figure 5: Error detection via reliability scores. (Left) ROC curve with AUC =0.801=0.801. Operating point at $\tau^{*}=0.445$: 72.5% true positive rate (error capture) at 27.0% false positive rate. (Right) Precision-recall curve. At operating point: 34.6% precision, 72.5% recall. Horizontal line: baseline 16.5% error rate. The $2.1\times$ improvement over random deferral demonstrates effective error identification.
AUC-ROC =0.801=0.801 indicates strong discriminative ability for separating errors from correct predictions. At threshold $\tau^{*}=0.445$: error capture (true positive rate) $72.5\%$; false positive rate $27.0\%$; precision $34.6\%$; improvement over random $72.5\%/34.5\%=2.1\times$. This instance-level discrimination is precisely what calibration alone (Step 1) cannot provide.
As a sanity check, we verified that simpler uncertainty heuristics—maximum predicted probability ($\max_{k}p_{k}$) and entropy ($\sum_{k}p_{k}\log p_{k}$)—achieve similar AUC ($\approx 0.80$) when used to rank samples by uncertainty. This confirms that the geometric reliability score successfully captures the uncertainty information present in calibrated probabilities. The advantage of the Fisher-Rao formulation lies not in empirical superiority on a single dataset, but in the theoretical guarantees of Theorem 2 (concentration bounds, sample efficiency) and the principled threshold selection via the $\alpha$-level constraint.
4.6 Neutral Zone Performance: Steps 1 and 2 Combined
The neutral zone mechanism combines both steps: calibrated probabilities (Step 1) feed into reliability scores (Step 2), which determine deferral decisions. Figure 6 presents the complete trade-off between automation and accuracy. The Pareto frontier reveals three regimes:
Figure 6: Pareto frontier: error rate vs. deferral rate. Operating point (red dot): 34.5% deferral, 6.9% error rate. Gray horizontal line: 16.8% baseline error rate. At 50% deferral, error rate reaches 3.9%.
High automation (0–30% deferral): Error rate drops rapidly from 16.8% to $\sim$8%—50% reduction while maintaining 70% automation.
Moderate automation (30–60%): Continued improvement with diminishing returns. Our operating point ($\tau^{*}=0.445$: 34.5% deferral, 6.9% error) lies at the curve’s knee. At 50% deferral, error rate reaches 3.9%.
Minimal automation (>>60%): Eliminating errors entirely requires $\sim$80% deferral—impractical for deployment.
4.6.1 Quantifying the Two-Stage Framework’s Value
Table 4 isolates contributions from each component.
Table 4: Decomposing the two-stage framework. Calibration alone (Step 1) provides marginal accuracy gains. Adding neutral zones (Step 2) achieves 59% error reduction for automated decisions.
Configuration Error Rate Automation Improvement
Uncalibrated CNN 16.8% 100%
+ Step 1 (calibration) 16.5% 100% $2\%$
+ Step 2 (neutral zone) 6.9% 65.5% $\mathbf{-59\%}$
The key insight: Calibration (Step 1) contributes modestly to accuracy but is essential as the foundation for Step 2. The reliability score $R$ is computed from calibrated probabilities; without proper calibration, the geometric distance to simplex vertices would not accurately reflect prediction reliability. The two steps are synergistic: neither alone achieves uncertainty-aware classification, but together they reduce automated decision errors by 59%.
4.6.2 Comparison to Alternative Paradigms
Random deferral: At 34.5% deferral, random selection yields no improvement (expected error rate remains 16.8%). Our method achieves 6.9%—a $2.4\times$ reduction.
Conformal prediction: Set-valued methods guarantee coverage but produce ambiguous outputs (e.g., $\{\text{Empty},\text{dsDNA}\}$) requiring human interpretation. Our approach provides crisp predictions with scalar reliability quantification, enabling straightforward integration into clinical workflows.
4.7 Sample Efficiency
Theorem 2 provides concentration bounds with explicit sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4\approx 0.229$ for $\lambda=1$. This has direct implications for validation set requirements. For reliability guarantee $\mathbb{P}(|R-\mathbb{E}[R]|>t)\leq\delta$ with $\delta=0.01$ and $t=0.1$:
$n_{\text{ours}}=\frac{0.229\times\log(200)}{0.02}\approx 61\text{ samples}$ (17)
A naive Lipschitz-diameter approach, treating the reliability function as having range $\lambda\pi$ without exploiting its exponential decay structure, yields sub-Gaussian parameter $\sigma^{2}_{\text{naive}}=(\lambda\pi)^{2}/4\approx 2.47$ for $\lambda=1$. This requires $n_{\text{naive}}\approx 654$ samples for the same guarantees—a 10.7-fold improvement from our tighter bound.  [^2] We emphasize that this comparison is against worst-case Lipschitz bounds that make no assumption about the reliability function’s structure. Practitioners who estimate the score’s range empirically from pilot data would obtain bounds similar to ours. The value of Theorem 2 lies in providing these bounds analytically, enabling sample size determination at the study design phase before validation data is collected.
Our empirical validation used $n=310$ samples, providing a safety margin of $310/61\approx 5\times$ above the theoretical minimum. This margin accommodates model misspecification and finite-sample effects not captured by asymptotic theory.
4.8 Robustness Analysis
4.8.1 Hyperparameter Sensitivity
Regularization ($\lambda_{1},\lambda_{2}$): Varying across $\{0.001,0.01,0.1\}$ yields error rates in $6.7\%,7.0\%]$—stable across two orders of magnitude.
Reliability parameter ($\lambda$): Testing $\lambda\in\{0.5,1.0,1.5,2.0\}$ produces identical AUC-ROC =0.80=0.80 across all values, indicating robustness to this parameter choice.
Threshold ($\tau^{*}$): Perturbations of $\pm 0.05$ yield error rates of 8.4% (at $\tau^{*}-0.05$) and 5.5% (at $\tau^{*}+0.05$), with corresponding deferral rates of 26.8% and 41.3%. The smooth Pareto frontier indicates stable operating regions.
4.8.2 Cross-Validation
Five-fold cross-validation yields: error rate $7.1\%\pm 5.5\%$; error capture $74.6\%\pm 13.4\%$; deferral rate $35.8\%\pm 7.0\%$. The relatively high variance reflects the small validation set size ($n=310$) and class imbalance (ssDNA comprises only 18.4% of samples).
4.8.3 Method Comparison
Table 5 compares geometric calibration against established post-hoc calibration methods: temperature scaling [1], Platt scaling [8], and isotonic regression [23]. For fair comparison, all methods use the same reliability score formula $)R=\exp(-\lambda\cdot d_{FR}(\mathbf{p}^{\text{cal}},\mathbf{e}_{\hat{j}}$)) applied to their respective calibrated probabilities, with thresholds learned to achieve 34.5% deferral.
Table 5: Post-hoc calibration method comparison at 34.5% deferral. All methods achieve similar error detection performance (AUC 0.78-0.80, capture 70-73%) when combined with reliability-score-based deferral. Geometric calibration achieves highest AUC but the advantage is modest. Platt scaling achieves best overall accuracy (84.5%) but slightly lower error detection. These results suggest the neutral zone mechanism itself—rather than the specific calibration method—drives operational gains.
Method Errors Accuracy ECE $\downarrow$ AUC $\uparrow$ Capture
Uncalibrated 52 83.2% 0.016 0.797 71.2%
Temperature 52 83.2% 0.022 0.797 73.1%
Platt (OvR) 48 84.5% 0.021 0.778 70.8%
Isotonic 50 83.9% 0.034 0.784 70.0%
Geometric 51 83.5% 0.022 0.801 72.5%
All methods achieve similar performance, with differences of only 2–3 percentage points across metrics. Platt scaling achieves the best overall accuracy (84.5%, 48 errors), while geometric calibration achieves the highest error detection AUC (0.801). These modest differences lead to an important conclusion about the paper’s contribution.
Interpretation. The operational gains demonstrated in Section 4.6—72.5% error capture at 34.5% deferral, yielding 6.9% automated error rate—arise primarily from the neutral zone mechanism itself, which can be applied to any calibration method. The specific contribution of geometric calibration is therefore not empirical superiority but rather:
  1. 1.
    A principled multi-class generalization of Platt scaling grounded in information geometry (Proposition 1), filling a theoretical gap in the calibration literature;
  2. 2.
    Formal consistency guarantees with explicit convergence rate (Theorem 1);
  3. 3.
    Concentration bounds with explicit sub-Gaussian parameters enabling sample size calculations before data collection (Theorem 2).
For applications where empirical performance is the sole criterion, simpler methods like temperature scaling may suffice. However, for settings requiring formal validation documentation—regulatory submissions, safety-critical deployments, or expensive validation studies where sample size must be justified a priori—the theoretical foundations of geometric calibration provide value that empirical benchmarks alone cannot capture.
5 Open Problems and Future Directions
This section outlines research directions extending the geometric calibration framework. We organize by theme, progressing from theoretical extensions to practical questions and connections with related methods.
5.1 Curvature-Based Error Bounds
The positive curvature $K=+1/4$ of the probability simplex suggests tighter error propagation than Euclidean space. Specifically, geodesic triangles on positively curved spaces have angle sums exceeding $\pi$, and comparison theorems (Toponogov, Rauch) bound distance distortions. We conjecture:
$d_{\text{FR}}(\mathbf{p},\mathbf{q})\leq 2\tanh^{-1}\left(\frac{\|\operatorname{ALR}(\mathbf{p})-\operatorname{ALR}(\mathbf{q})\|_{2}}{C}\right)$ (18)
for some constant $C>0$, with the sublinear $\tanh^{-1}$ growth suggesting tighter bounds than linear Euclidean propagation. The intuition is that the ALR transformation “stretches” distances near the simplex boundary, while the spherical geometry “compresses” them, potentially yielding favorable cancellation.
5.2 Distribution Shift Robustness
Under covariate shift $Q\ll P$ with bounded likelihood ratio $dQ/dP\leq M$, we aim to prove:
$\left|\operatorname{\mathbb{E}}_{Q}[d_{\text{FR}}]-\operatorname{\mathbb{E}}_{P}[d_{\text{FR}}]\right|\leq\sqrt{(M-1)^{2}+2M\cdot D_{KL}(Q\|P)}\cdot\pi$ (19)
using the bounded diameter $\operatorname{diam}(\Delta^{c-1})=\pi$. This would provide guarantees for calibration transfer across related domains (e.g., training on one imaging modality, deploying on another).
5.3 Geometry vs. Temperature Scaling: Phase Transitions
Temperature scaling uses one parameter; geometric calibration uses $(c-1)^{2}+(c-1)$. We hypothesize a phase transition: geometric methods dominate when miscalibration is anisotropic, quantified by effective rank:
$r_{\text{eff}}=\frac{\operatorname{tr}(\Sigma)}{\|\Sigma\|_{2}},\quad\Sigma=\operatorname{Cov}[\operatorname{ALR}(\mathbf{p}^{\text{CNN}})].$ (20)
When $r_{\text{eff}}\approx 1$, miscalibration is nearly isotropic (all directions equally affected), and temperature scaling suffices. When $r_{\text{eff}}\gg 1$, different ALR directions require different corrections, favoring the full geometric approach. Characterizing the critical threshold $r^{*}_{\text{eff}}(c)$ as a function of number of classes remains open.
5.4 Neutral Zone Size Reduction
Does calibration reduce neutral zone size exponentially? Via Fano’s inequality, we conjecture:
$|NZ^{\text{cal}}|/|NZ^{\text{uncal}}|\stackrel{{\scriptstyle?}}{{\approx}}\exp(-\theta\cdot\Delta\text{ECE})$ (21)
for some $\theta>0$ depending on the class structure. The intuition is that improved calibration concentrates predictions away from decision boundaries, reducing the proportion requiring deferral. Empirical validation across diverse datasets would clarify the functional form.
5.5 Extensions to r-Power Methodology
The following directions extend the r-power framework [12] and will be developed in subsequent work.
5.5.1 Multi-Class Extension
Extension beyond binary classification requires testing ordered hypotheses via Union-Intersection principles. Specifically, we test:
$H_{0}^{(j)}:p_{j}\leq\max_{k\neq j}p_{k}\quad\text{vs.}\quad H_{1}^{(j)}:p_{j}>\max_{k\neq j}p_{k}$ (22)
for the predicted class $j=\hat{j}$. Our current approach uses step-down methods, adjusting for conservativeness through controlled false negative rates. The key challenge is maintaining power while controlling familywise error across the c(c1)/2c(c-1)/2 pairwise comparisons implicit in the multi-class setting.
5.5.2 ADEBO Adaptation for Bounded Support
The ADEBO estimator [24] requires approximate normality. For probability vectors with bounded support $\epsilon,1-\epsilon]^{c}\cap\Delta^{c-1}$, the logit transformation induces skewness near boundaries. Deriving moment conditions for bounded support—particularly higher-order corrections to the normal approximation—would extend ADEBO’s applicability to calibrated probability outputs.
6 Conclusion
We have developed a geometric framework for uncertainty-aware classification that bridges information geometry and statistical learning theory. The framework comprises two stages: geometric calibration of neural network probability outputs via ALR transformations, followed by reliability scoring and neutral zone construction for principled deferral of uncertain predictions.
Main contributions.
Our central theoretical result is Proposition 1, which establishes geometric calibration as the natural multi-class generalization of Platt scaling—a widely-used binary method that previously lacked principled extension to $c>2$ classes. This fills a gap in the calibration literature by grounding multi-class calibration in the intrinsic geometry of the probability simplex.
Supporting this framework, Theorem 1 provides consistency guarantees at rate $O_{p}(n^{-1/2})$ via M-estimation theory, while Theorem 2 yields concentration bounds for reliability scores with explicit sub-Gaussian parameters. These bounds enable practitioners to determine required validation set sizes before data collection—a practical benefit for applications where labeled validation data is expensive.
Empirical findings.
Validation on AAV classification demonstrates that the two-stage framework achieves substantial error reduction: 72.5% error capture at 34.5% deferral rate, reducing automated decision errors from 16.8% to 6.9%. Importantly, our comparison with alternative calibration methods (Table 5) reveals that these operational gains arise primarily from the neutral zone mechanism itself, which is applicable to any well-calibrated probability output. Geometric calibration achieves competitive empirical performance while providing formal guarantees that simpler methods lack.
Limitations.
Several limitations warrant acknowledgment. First, empirical validation is confined to a single application domain with $c=3$ classes; the framework’s behavior in high-dimensional settings ($c\gg 3$) and under distribution shift remains to be demonstrated. Second, the evidence for faster-than-$n^{-1/2}$ convergence (Remark 11) is suggestive but inconclusive, and Conjecture 1 on Neyman-Pearson optimality lacks rigorous proof. Third, the cross-validation variance (Section 4.8) indicates that operating point performance exhibits meaningful uncertainty at $n=310$.
Future directions.
The open problems in Section 5 outline extensions of this work. Most pressing are: (i) rigorous investigation of Conjecture 1, potentially via explicit characterization of the likelihood ratio under Dirichlet perturbations; (ii) distribution shift robustness bounds for calibration transfer across related domains; and (iii) integration with the r-power methodology [12] for distribution-free neutral zones in multi-class settings. The connection between Fisher-Rao geometry and optimal decision theory suggested by Remark 6 merits deeper exploration.
Broader context.
As AI systems increasingly inform high-stakes decisions in medicine, criminal justice, and autonomous systems, principled uncertainty quantification becomes essential. This work contributes formal foundations for one approach to this challenge. The theoretical guarantees established here—consistency rates, concentration bounds, and the principled connection to Platt scaling—may prove particularly valuable in applications requiring rigorous validation documentation, where empirical performance benchmarks alone are insufficient.
Appendix A Proofs
A.1 Proof of Theorem 1
We prove consistency $\hat{T}_{n}\xrightarrow{p}T^{*}$ and the rate $\|\hat{T}_{n}-T^{*}\|=O_{p}(n^{-1/2})$ using the M-estimation framework of van der Vaart [20].
Preliminary: M-Estimation Setup
Our estimator minimizes an empirical loss:
$\hat{T}_{n}=\operatorname*{arg\,min}_{T\in\mathcal{T}}\mathcal{L}_{n}(T),\quad\mathcal{L}_{n}(T)=\frac{1}{n}\sum_{i=1}^{n}\ell(T;\mathbf{p}_{i},y_{i})+\lambda_{1}\|A-I\|_{F}^{2}+\lambda_{2}\|\mathbf{b}\|^{2}$ (23)
where the per-sample loss is $\ell(T;\mathbf{p},\mathbf{y})=-\sum_{j=1}^{c}y_{j}\log p_{j}^{\text{cal}}(T)$.
The population version is:
$T^{*}=\operatorname*{arg\,min}_{T\in\mathcal{T}}\mathcal{L}^{*}(T),\quad\mathcal{L}^{*}(T)=\mathbb{E}[\ell(T;\mathbf{p},\mathbf{y})]+\lambda_{1}\|A-I\|_{F}^{2}+\lambda_{2}\|\mathbf{b}\|^{2}$ (24)
Remark 14 (Van der Vaart’s Convention).
Van der Vaart’s theorems are stated for maximizing a criterion $M_{n}(\theta)$. To apply them to our minimization problem, we work with $M_{n}(T)=-\mathcal{L}_{n}(T)$ and $M^{*}(T)=-\mathcal{L}^{*}(T)$. All conditions below are verified for this flipped objective.
Part I: Consistency
Van der Vaart, Theorem 5.7: Let $\hat{\theta}_{n}$ maximize $M_{n}(\theta)$ over a set $\Theta$. If:
  1. (i)
    $\Theta$ is compact,
  2. (ii)
    $\theta\mapsto M(\theta)$ is continuous,
  3. (iii)
    $\sup_{\theta\in\Theta}|M_{n}(\theta)-M(\theta)|\xrightarrow{p}0$,
  4. (iv)
    For all $\varepsilon>0$: $\sup_{\theta:d(\theta,\theta_{0})\geq\varepsilon}M(\theta)<M(\theta_{0})$ (well-separation),
then $\hat{\theta}_{n}\xrightarrow{p}\theta_{0}$.
We verify each condition for $M_{n}=-\mathcal{L}_{n}$ and $M^{*}=$-$\mathcal{L}^{*}$.
Condition (i): Compactness of $\mathcal{T}$.
Define the parameter space:
$\mathcal{T}=\{(A,\mathbf{b}):\|A\|_{F}\leq M_{A},\|\mathbf{b}\|\leq M_{b},A\succeq\delta I,\operatorname{tr}(A)=c-1\}$ (25)
where $M_{A},M_{b},\delta>0$ are determined by the regularization strength. The lower bound $\delta>0$ on eigenvalues ensures $\mathcal{T}$ is closed (and hence compact, being bounded). In practice, the regularization term $\lambda_{1}\|A-I\|_{F}^{2}$ keeps $A$ well-conditioned: for $\|A-I\|_{F}\leq M_{A}$ with $M_{A}<1$, we have $\lambda_{\min}(A)\geq 1-M_{A}>0$.
Proof of compactness.
We show $\mathcal{T}$ is closed and bounded in $\mathbb{R}^{(c-1)^{2}+(c-1)}$.
Closed: $\mathcal{T}$ is the intersection of:
  • $\{(A,\mathbf{b}):\|A\|_{F}\leq M_{A}\}$ is closed (continuous norm, inverse image of $0,M_{A}]$)
  • $\{(A,\mathbf{b}):\|\mathbf{b}\|\leq M_{b}\}$ is closed
  • $\{(A,\mathbf{b}):A\succeq\delta I\}$ is closed (smallest eigenvalue is continuous, inverse image of $\delta,\infty)$)
  • $\{(A,\mathbf{b}):\operatorname{tr}(A)=c-1\}$ is closed (continuous trace, inverse image of singleton)
Finite intersection of closed sets is closed.
Bounded: $\|A\|_{F}\leq M_{A}$ and $\|\mathbf{b}\|\leq M_{b}$ directly.
By the Heine-Borel theorem, $\mathcal{T}$ is compact. ∎
Condition (ii): Continuity of $\mathcal{L}^{*}$.
The regularization terms $\lambda_{1}\|A-I\|_{F}^{2}+\lambda_{2}\|\mathbf{b}\|^{2}$ are continuous. For the expectation term, we show $T\mapsto\mathbb{E}[\ell(T;\mathbf{p},\mathbf{y})]$ is continuous.
Proof of continuity.
Step 1: Pointwise continuity. For each fixed $(\mathbf{p},\mathbf{y})$, the map $T\mapsto\ell(T;\mathbf{p},\mathbf{y})$ is continuous because:
$\ell(T;\mathbf{p},\mathbf{y})=-\sum_{j}y_{j}\log p_{j}^{\text{cal}}(T),\quad p_{j}^{\text{cal}}(T)=\text{softmax}_{j}(A\cdot\text{ALR}(\mathbf{p})+\mathbf{b})$ (26)
This is a composition of:
  • $(A,\mathbf{b})\mapsto A\mathbf{z}+\mathbf{b}$ (linear, continuous)
  • $\mathbf{z}^{\text{cal}}\mapsto\text{softmax}(\mathbf{z}^{\text{cal}})$ (smooth, $C^{\infty}$)
  • $\mathbf{p}^{\text{cal}}\mapsto-\sum_{j}y_{j}\log p_{j}^{\text{cal}}$ (continuous when $p_{j}>0$)
Since $p_{j}^{\text{cal}}\geq\epsilon^{\prime}>0$ for all $T\in\mathcal{T}$ (Lemma 1), the logarithm is well-defined and continuous.
Step 2: Dominated convergence. By Lemma 1, there exists a constant $M<\infty$ such that $|\ell(T;\mathbf{p},\mathbf{y})|\leq M$ for all $T\in\mathcal{T}$, with $\mathbb{E}[M]=M<\infty$.
For any sequence $T_{n}\to T_{0}$ in $\mathcal{T}$:
$\lim_{n\to\infty}\mathbb{E}[\ell(T_{n};\mathbf{p},\mathbf{y})]=\mathbb{E}\left[\lim_{n\to\infty}\ell(T_{n};\mathbf{p},\mathbf{y})\right]=\mathbb{E}[\ell(T_{0};\mathbf{p},\mathbf{y})]$ (27)
Therefore $\mathcal{L}^{*}$ is continuous. ∎
Condition (iii): Uniform Convergence (Glivenko-Cantelli).
We must show $\sup_{T\in\mathcal{T}}|\mathcal{L}_{n}(T)-\mathcal{L}^{*}(T)|\xrightarrow{p}0$.
Since the regularization terms are deterministic, it suffices to prove:
$\sup_{T\in\mathcal{T}}\left|\frac{1}{n}\sum_{i=1}^{n}\ell(T;\mathbf{p}_{i},y_{i})-\mathbb{E}[\ell(T;\mathbf{p},\mathbf{y})]\right|\xrightarrow{p}0$ (28)
Van der Vaart, Example 19.8 (Pointwise Compact Class): Let $\mathcal{F}=\{f_{\theta}:\theta\in\Theta\}$ be a parametric class indexed by a compact metric space $\Theta$ such that $\theta\mapsto f_{\theta}(x)$ is continuous for every $x$, with integrable envelope function $F$. Then the $L_{1}$-bracketing numbers of $\mathcal{F}$ are finite and hence $\mathcal{F}$ is Glivenko-Cantelli.
Verification of Glivenko-Cantelli conditions.
  1. 1.
    $\mathcal{T}$ is a compact metric space (Condition (i)) $\clubsuit$
  2. 2.
    For each fixed $(\mathbf{p},\mathbf{y})$, the map $T\mapsto\ell(T;\mathbf{p},\mathbf{y})$ is continuous (Condition (ii)) $\clubsuit$
  3. 3.
    Integrable envelope: $|\ell(T;\mathbf{p},\mathbf{y})|\leq M$ for all $T\in\mathcal{T}$ with $\mathbb{E}[M]=M<\infty$ (Lemma 1) $\clubsuit$
By Example 19.8, the function class is Glivenko-Cantelli. ∎
Lemma 1 (Bounded Loss).
Under the assumption $p_{j}^{\text{CNN}}\geq\epsilon>0$ for all $j=1,\ldots,c$, there exists a constant $M<\infty$ (depending only on $\epsilon,M_{A},M_{b},c$) such that:
$\sup_{T\in\mathcal{T}}|\ell(T;\mathbf{p},\mathbf{y})|\leq M$ (29)
Proof.
We establish the bound through a chain of four inequalities.
Step (a): Bounded ALR inputs. Since $p_{j}^{\text{CNN}}\geq\epsilon$ and $\sum_{j}p_{j}=1$:
$p_{j}^{\text{CNN}}\leq 1-\sum_{k\neq j}p_{k}^{\text{CNN}}\leq 1-(c-1)\epsilon$ (30)
The ALR coordinates satisfy:
$|z_{k}|=\left|\log\frac{p_{k}}{p_{c}}\right|\leq\max\left\{\log\frac{1-(c-1)\epsilon}{\epsilon},\log\frac{\epsilon}{1-(c-1)\epsilon}\right\}=:B_{z}<\infty$ (31)
Therefore $\|\mathbf{z}\|_{2}\leq B_{z}\sqrt{c-1}$.
Step (b): Bounded calibrated logits. For any $T=(A,\mathbf{b})\in\mathcal{T}$:
$\|\mathbf{z}^{\text{cal}}\|_{2}=\|A\mathbf{z}+\mathbf{b}\|_{2}\leq\|A\|_{op}\|\mathbf{z}\|_{2}+\|\mathbf{b}\|_{2}\leq M_{A}\cdot B_{z}\sqrt{c-1}+M_{b}=:B_{\text{cal}}$ (32)
where we used $\|A\|_{op}\leq\|A\|_{F}\leq M_{A}$. Thus $\|\mathbf{z}^{\text{cal}}\|_{\infty}\leq B_{\text{cal}}$.
Step (c): Lower bound on calibrated probabilities. For softmax with bounded inputs:
$p_{j}^{\text{cal}}=\frac{e^{z_{j}^{\text{cal}}}}{\sum_{k=1}^{c}e^{z_{k}^{\text{cal}}}}\geq\frac{e^{-B_{\text{cal}}}}{c\cdot e^{B_{\text{cal}}}}=\frac{1}{c}e^{-2B_{\text{cal}}}=:\epsilon^{\prime}>0$ (33)
Step (d): Bounded log-loss. Since $\mathbf{y}$ is one-hot encoded ($\sum_{j}y_{j}=1$, exactly one $y_{j}=1$):
$|\ell(T;\mathbf{p},\mathbf{y})|=\left|-\sum_{j}y_{j}\log p_{j}^{\text{cal}}\right|=-\log p_{\hat{y}}^{\text{cal}}\leq-\log(\epsilon^{\prime})=\log(c)+2B_{\text{cal}}=:M$ (34)
where $\hat{y}$ is the true class. Critically, $M$ depends only on constants $(\epsilon,M_{A},M_{b},c)$, not on the random data. ∎
Condition (iv): Well-Separation (Identifiability).
For the flipped objective $M^{*}=$-$\mathcal{L}^{*}$, van der Vaart’s condition requires:
$\sup_{T:\|T-T^{*}\|\geq\varepsilon}M^{*}(T)<M^{*}(T^{*})$ (35)
Equivalently (multiplying by 1-1 and reversing inequality):
$\inf_{T:\|T-T^{*}\|\geq\varepsilon}\mathcal{L}^{*}(T)>\mathcal{L}^{*}(T^{*})$ (36)
This follows from strong convexity.
Claim 1.
$\mathcal{L}^{*}$ is strongly convex with modulus $\mu=\min\{2\lambda_{1},2\lambda_{2}\}>0$.
Proof.
The Hessian decomposes as:
$\nabla^{2}\mathcal{L}^{*}(T)=\mathbb{E}[\nabla^{2}_{T}\ell(T;\mathbf{p},\mathbf{y})]+2\lambda_{1}I_{(c-1)^{2}}+2\lambda_{2}I_{c-1}$ (37)
To connect Lemma 2 to the parameter Hessian, we use the chain rule. Let $J_{T}=\partial\mathbf{z}^{\text{cal}}/\partial T$ denote the Jacobian of the map $T\mapsto A\cdot\text{ALR}(\mathbf{p})+\mathbf{b}$. The Hessian of the per-sample loss decomposes as:
$\nabla^{2}_{T}\ell(T;\mathbf{p},y)=J_{T}^{\top}\Sigma_{\mathbf{p}^{\text{cal}}}J_{T}$ (38)
where $\Sigma_{\mathbf{p}}=\text{diag}(\mathbf{p})-\mathbf{p}\mathbf{p}^{\top}$ is the softmax Hessian. By Lemma 2, $\Sigma_{\mathbf{p}^{\text{cal}}}\succeq\epsilon^{2}I_{c-1}$, which implies:
$\mathbb{E}[\nabla^{2}_{T}\ell]=\mathbb{E}[J_{T}^{\top}\Sigma_{\mathbf{p}^{\text{cal}}}J_{T}]\succeq 0$ (39)
The data term is positive semi-definite but not uniformly positive definite, since $J_{T}\in\mathbb{R}^{(c-1)\times[(c-1)^{2}+(c-1)]}$ has rank at most c1c-1, leaving a non-trivial nullspace. The regularization terms provide the requisite uniform curvature:
$\nabla^{2}\mathcal{L}^{*}(T)\succeq\mu I,\quad\mu=\min\{2\lambda_{1},2\lambda_{2}\}>0$ (40)
Strong convexity implies: for any $T\neq T^{*}$,
$\mathcal{L}^{*}(T)\geq\mathcal{L}^{*}(T^{*})+\nabla\mathcal{L}^{*}(T^{*})^{\top}(T-T^{*})+\frac{\mu}{2}\|T-T^{*}\|^{2}$ (41)
At the minimizer, $\nabla\mathcal{L}^{*}(T^{*})=0$, so:
$\mathcal{L}^{*}(T)\geq\mathcal{L}^{*}(T^{*})+\frac{\mu}{2}\|T-T^{*}\|^{2}$ (42)
For all $T$ with $\|T-T^{*}\|\geq\varepsilon$:
$\inf_{\|T-T^{*}\|\geq\varepsilon}\mathcal{L}^{*}(T)\geq\mathcal{L}^{*}(T^{*})+\frac{\mu\varepsilon^{2}}{2}>\mathcal{L}^{*}(T^{*})$ (43)
Strong convexity also ensures $T^{*}$ is the unique minimizer. Existence follows from compactness of $\mathcal{T}$ and continuity of $\mathcal{L}^{*}$ (Weierstrass extreme value theorem).
Conclusion.
All four conditions of van der Vaart Theorem 5.7 are satisfied. Therefore:
$\boxed{\hat{T}_{n}\xrightarrow{p}T^{*}}$ (44)
Part II: Rate
Van der Vaart, Theorem 5.23: Let $\hat{\theta}_{n}$ maximize $M_{n}(\theta)$. If in addition to Theorem 5.7:
  1. (i)
    $\theta\mapsto M_{n}(\theta)$ is twice continuously differentiable,
  2. (ii)
    $\theta_{0}\in\text{interior}(\Theta)$,
  3. (iii)
    The map $\theta\mapsto PM_{\theta}$ admits a second-order Taylor expansion at $\theta_{0}$ with non-singular Hessian $V_{\theta_{0}}$,
  4. (iv)
    $\mathbb{P}_{n}M_{\hat{\theta}_{n}}\geq\sup_{\theta}\mathbb{P}_{n}M_{\theta}-o_{P}(n^{-1})$,
  5. (v)
    Stochastic equicontinuity holds,
then:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\xrightarrow{d}N(0,V_{\theta_{0}}^{-1}\text{Var}(\dot{M}_{\theta_{0}})V_{\theta_{0}}^{-1})$ (45)
Condition (i): Twice Continuous Differentiability.
The loss $\ell(T;\mathbf{p},\mathbf{y})=-\sum_{j}y_{j}\log p_{j}^{\text{cal}}(T)$ is $C^{\infty}$ in $T$ since softmax is infinitely differentiable and $p_{j}^{\text{cal}}\geq\epsilon^{\prime}>0$ ensures logarithm is smooth. Therefore $\mathcal{L}_{n}$ and $\mathcal{L}^{*}$ are twice continuously differentiable. $\clubsuit$
Condition (ii): Interior Solution.
For sufficiently large $M_{A}$ and $M_{b}$ (chosen based on the data-generating process), the minimizer $T^{*}$ satisfies $\|A^{*}\|_{F}<M_{A}$ and $\|\mathbf{b}^{*}\|<M_{b}$ strictly. The regularization terms penalize large parameters, ensuring the solution lies strictly in the interior. $\clubsuit$
Condition (iii): Non-Singular Hessian.
We have $H:=\nabla^{2}\mathcal{L}^{*}(T^{*})\succeq\mu I$ with $\mu>0$ from strong convexity (Part I). $\clubsuit$
Condition (iv): Optimality.
Our $\hat{T}_{n}$ is the exact minimizer: $\mathcal{L}_{n}(\hat{T}_{n})=\min_{T}\mathcal{L}_{n}(T)$. Flipping signs: $M_{n}(\hat{T}_{n})=\max_{T}M_{n}(T)$, which satisfies the condition with $o_{P}(n^{-1})=0$. $\clubsuit$
Condition (v): Stochastic Equicontinuity.
The gradient satisfies $\nabla_{T}\ell(T;\mathbf{p},\mathbf{y})=J_{T}^{\top}(\mathbf{p}^{\text{cal}}-\mathbf{y})$ where $J_{T}$ is the Jacobian. On compact $\mathcal{T}$, both $J_{T}$ and $\mathbf{p}^{\text{cal}}-\mathbf{y}$ are bounded, giving $\|\nabla_{T}\ell\|_{2}\leq C<\infty$. This implies finite variance: $\text{Var}(\nabla_{T}\ell(T^{*};\mathbf{p},\mathbf{y}))\leq C^{2}<\infty$. $\clubsuit$
Lemma 2 (Softmax Hessian Eigenvalue).
For $\mathbf{p}\in\Delta^{c-1}$ with $p_{j}\geq\epsilon$ for all $j$, the softmax Hessian satisfies:
$\lambda_{\min}(\Sigma_{\mathbf{p}})\geq\epsilon^{2},\quad\Sigma_{\mathbf{p}}=\text{diag}(\mathbf{p})-\mathbf{p}\mathbf{p}^{\top}$ (46)
Proof.
Work with the $(c-1)\times(c-1)$ matrix $H=\text{diag}(p_{1},\ldots,p_{c-1})-\mathbf{p}_{1:c-1}\mathbf{p}_{1:c-1}^{\top}$. For any unit vector $\mathbf{v}\in\mathbb{R}^{c-1}$ with $\|\mathbf{v}\|_{2}=1$:
$\mathbf{v}^{\top}H\mathbf{v}$ $=\sum_{i=1}^{c-1}p_{i}v_{i}^{2}-\left(\sum_{i=1}^{c-1}p_{i}v_{i}\right)^{2}$ (47)
$\geq\sum_{i=1}^{c-1}p_{i}v_{i}^{2}-(1-p_{c})\sum_{i=1}^{c-1}p_{i}v_{i}^{2}\quad\text{(Cauchy-Schwarz: $(\sum a_{i}b_{i})^{2}\leq(\sum a_{i})(\sum a_{i}b_{i}^{2})$)}$ (48)
$=p_{c}\sum_{i=1}^{c-1}p_{i}v_{i}^{2}$ (49)
$\geq p_{c}\cdot\min_{1\leq i\leq c-1}p_{i}\cdot\underbrace{\sum_{i=1}^{c-1}v_{i}^{2}}_{=\|\mathbf{v}\|_{2}^{2}=1}$ (50)
$\geq\epsilon\cdot\epsilon=\epsilon^{2}$ (51)
Conclusion.
By van der Vaart Theorem 5.23:
$\sqrt{n}(\hat{T}_{n}-T^{*})\xrightarrow{d}N(0,H^{-1}VH^{-1})$ (52)
where $H=\nabla^{2}\mathcal{L}^{*}(T^{*})$ and $V=\text{Var}(\nabla\ell(T^{*};\mathbf{p},\mathbf{y}))$.
This implies:
$\boxed{\|\hat{T}_{n}-T^{*}\|=O_{p}(n^{-1/2})}$ (53)
A.2 Proof of Theorem 2
We prove concentration of the reliability score $)R(\mathbf{p})=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{e}_{\hat{j}}$)) where $\hat{j}=\operatorname*{arg\,max}_{k}p_{k}$.
Remark 15 (Proof Strategy).
The reliability function involves an $\operatorname*{arg\,max}$ operation, which could introduce discontinuities at decision boundaries where $p_{i}=p_{j}$ for some $i\neq j$. We first prove $R$ is continuous everywhere (Steps 1–3), then use this continuity with the bounded range property to apply Hoeffding’s inequality (Steps 4–6).
Step 1: Decision Region Decomposition
The simplex $\Delta^{c-1}$ partitions into $c$ decision regions:
$\Delta_{j}=\{\mathbf{p}\in\Delta^{c-1}:p_{j}=\max_{k}p_{k}\},\quad j=1,\ldots,c$ (54)
Within each $\Delta_{j}$, the predicted class is fixed at $\hat{j}=j$, so the map $R_{j}(\mathbf{p})=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j}))$ is smooth (composition of smooth functions: Fisher-Rao distance and exponential).
The decision boundaries $\mathcal{B}=\{\mathbf{p}:p_{i}=p_{j}\text{ for some }i\neq j\}$ form a set of measure zero in $\Delta^{c-1}$ (union of lower-dimensional submanifolds).
Step 2: Lipschitz Property Within Regions
We establish Lipschitz continuity within each decision region, which will be used to prove global continuity in Step 3.
Claim 2.
$R_{j}$ is $\lambda$-Lipschitz on $\Delta_{j}$.
Proof.
The Fisher-Rao distance satisfies the triangle inequality. For any $\mathbf{p},\mathbf{q}\in\Delta^{c-1}$:
$|d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j})-d_{\text{FR}}(\mathbf{q},\mathbf{e}_{j})|\leq d_{\text{FR}}(\mathbf{p},\mathbf{q})$ (55)
For the exponential function $f(d)=e^{-\lambda d}$, we have $|f^{\prime}(d)|=\lambda e^{-\lambda d}\leq\lambda$ for all $d\geq 0$. By the mean value theorem and chain rule, for $\mathbf{p},\mathbf{q}\in\Delta_{j}$:
$|R_{j}(\mathbf{p})-R_{j}(\mathbf{q})|=|f(d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j}))-f(d_{\text{FR}}(\mathbf{q},\mathbf{e}_{j}))|\leq\lambda\cdot|d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j})-d_{\text{FR}}(\mathbf{q},\mathbf{e}_{j})|$ (56)
Applying the triangle inequality:
$|R_{j}(\mathbf{p})-R_{j}(\mathbf{q})|\leq\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{q})$ (57)
Step 3: Continuity Across Boundaries
Remark 16 (Key Insight).
Although the $\operatorname*{arg\,max}$ operation switches between regions at boundaries, the function values match exactly at these boundaries.
Consider a boundary point between $\Delta_{i}$ and $\Delta_{j}$ where $p_{i}=p_{j}$ (both are maximal). The Fisher-Rao distance to a vertex $\mathbf{e}_{k}$ is:
$d_{\text{FR}}(\mathbf{p},\mathbf{e}_{k})=2\arccos\left(\sum_{j=1}^{c}\sqrt{p_{j}\cdot\delta_{jk}}\right)=2\arccos(\sqrt{p_{k}})$ (58)
When $p_{i}=p_{j}$ at a boundary:
$d_{\text{FR}}(\mathbf{p},\mathbf{e}_{i})=2\arccos(\sqrt{p_{i}})=2\arccos(\sqrt{p_{j}})=d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j})$ (59)
Therefore:
$R_{i}(\mathbf{p})=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{e}_{i}))=\exp(-\lambda\cdot d_{\text{FR}}(\mathbf{p},\mathbf{e}_{j}))=R_{j}(\mathbf{p})$ (60)
The function values agree exactly at boundaries.
Conclusion.
$R(\mathbf{p})$ is continuous on all of $\Delta^{c-1}$. The $\operatorname*{arg\,max}$ operation introduces no discontinuity because the competing regions have identical function values where they meet. Combined with Lipschitz continuity within each region (Step 2), this establishes global continuity.
Step 4: Bounded Range
Lemma 3 (Diameter of Probability Simplex).
$\operatorname{diam}(\Delta^{c-1})=\pi$ under the Fisher-Rao metric.
Proof.
The maximum distance occurs between distinct vertices:
$d_{\text{FR}}(\mathbf{e}_{i},\mathbf{e}_{j})=2\arccos\left(\sum_{k=1}^{c}\sqrt{\delta_{ik}\cdot\delta_{jk}}\right)=2\arccos(0)=\pi\quad\text{for }i\neq j$ (61)
Since $d_{\text{FR}}(\mathbf{p},\mathbf{e}_{\hat{j}})\in[0,\pi]$, the reliability score has range:
$R\in[\exp(-\lambda\pi),1]$ (62)
  • Maximum: $R=1$ when $\mathbf{p}=\mathbf{e}_{\hat{j}}$ (perfect confidence in predicted class)
  • Minimum: $R=\exp(-\lambda\pi)$ when $\mathbf{p}=\mathbf{e}_{k}$ for $k\neq\hat{j}$ (maximum confusion)
The effective range is:
$\Delta R=1-\exp(-\lambda\pi)$ (63)
For common values of $\lambda$:
$\lambda$ $\exp(-\lambda\pi)$ Range $\Delta R$ Interpretation
0.5 0.208 0.792 Low sensitivity
1.0 0.043 0.957 Moderate sensitivity
2.0 0.002 0.998 High sensitivity
Step 5: Concentration via Hoeffding
Hoeffding’s Inequality [25]: Let $X_{1},\ldots,X_{n}$ be independent random variables with $X_{i}\in[a_{i},b_{i}]$ almost surely. Then for $\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$:
$\mathbb{P}(|\bar{X}-\mathbb{E}[\bar{X}]|>t)\leq 2\exp\left(-\frac{2n^{2}t^{2}}{\sum_{i=1}^{n}(b_{i}-a_{i})^{2}}\right)$ (64)
Application to our setting.
For i.i.d. observations with $R_{i}\in[e^{-\lambda\pi},1]$ having common range $\Delta R=1-e^{-\lambda\pi}$, Hoeffding’s inequality gives:
$\mathbb{P}(|\bar{R}_{n}-\mathbb{E}[R]|>t)\leq 2\exp\left(-\frac{2n^{2}t^{2}}{n(\Delta R)^{2}}\right)=2\exp\left(-\frac{2nt^{2}}{(\Delta R)^{2}}\right)$ (65)
For a single random draw ($n=1$), this yields:
$\boxed{\mathbb{P}(|R-\mathbb{E}[R]|>t)\leq 2\exp\left(-\frac{2t^{2}}{(1-e^{-\lambda\pi})^{2}}\right)}$ (66)
This proves part (i) of Theorem 2. ∎
Step 6: Sub-Gaussian Parameter
By Hoeffding’s lemma, any random variable $X$ with range $[a,b]$ satisfies:
$\mathbb{E}[e^{s(X-\mathbb{E}[X])}]\leq\exp\left(\frac{s^{2}(b-a)^{2}}{8}\right)$ (67)
This is the definition of a sub-Gaussian random variable with parameter $\sigma^{2}=(b-a)^{2}/4$.
For $R\in[e^{-\lambda\pi},1]$:
$\boxed{\sigma^{2}=\frac{(1-e^{-\lambda\pi})^{2}}{4}}$ (68)
This proves part (ii) of Theorem 2.
Remark 17 (Numerical values).
  • For $\lambda=1$: $\sigma^{2}=(1-e^{-\pi})^{2}/4\approx 0.916/4\approx 0.229$, giving $\sigma\approx 0.48$
  • For $\lambda=2$: $\sigma^{2}\approx 0.250$, giving $\sigma\approx 0.50$
This proves part (iii): for $\lambda=1$, the concentration bound becomes:
$\mathbb{P}(|R-\mathbb{E}[R]|>t)\leq 2\exp(-2.18\,t^{2})$ (69)
Comparison of Bounds
Alternative concentration bounds could be derived using Lipschitz properties. We compare approaches:
Approach Variance Bound For $\lambda=1$ Method
Lipschitz + diameter $\lambda^{2}\pi^{2}=9.87$ $2\exp(-0.20\,t^{2})$ McDiarmid
Hoeffding on range $(1-e^{-\pi})^{2}=0.916$ $\mathbf{2\exp(-2.18\,t^{2})}$ Hoeffding
The Hoeffding bound using the explicit range $e^{-\lambda\pi},1]$ is approximately $11\times$ tighter than a naive Lipschitz-diameter bound for typical $\lambda\approx 1$.
Remark 18 (Why the improvement?).
The Lipschitz approach uses diameter $\pi$ as the worst-case distance, but the reliability function decays exponentially with distance, making the effective range $1-e^{-\lambda\pi}\ll\lambda\pi$ for $\lambda\geq 1$.
Conclusion.
We have proven all three parts of Theorem 2:
  1. (i)
    The tail bound via Hoeffding’s inequality (Step 5)
  2. (ii)
    The sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$ (Step 6)
  3. (iii)
    The explicit constant $2.18$ for $\lambda=1$ (Step 6)
The key technical insight is that the reliability score $R(\mathbf{p})$ is continuous everywhere on $\Delta^{c-1}$ despite the $\operatorname*{arg\,max}$ operation in its definition (Step 3). This continuity, combined with the bounded range from the simplex diameter (Step 4), enables application of Hoeffding’s inequality to obtain a tight concentration bound.

### v2

@@
 Abstract
-Modern artificial intelligence systems make critical decisions yet often fail silently when uncertain.
-We develop a geometric framework for post-hoc calibration of neural network probability outputs, treating probability vectors as points on the (c1)(c-1)-dimensional probability simplex equipped with the Fisher–Rao metric.
-Our approach yields Additive Log-Ratio (ALR) calibration maps that reduce exactly to Platt scaling for binary problems (Proposition 1) while extending naturally to multi-class settings—providing a principled generalization that existing methods lack.
-Complementing calibration, we define geometric reliability scores based on Fisher–Rao distance and construct neutral zones for principled deferral of uncertain predictions.
-Theoretical contributions include: (i) consistency of the calibration estimator at rate $O_{p}(n^{-1/2})$ via M-estimation theory (Theorem 1), and (ii) tight concentration bounds for reliability scores with explicit sub-Gaussian parameters enabling sample size calculations for validation set design (Theorem 2).
+Modern artificial intelligence systems make critical decisions yet often fail silently when uncertain—even well-calibrated models provide no mechanism to identify which specific predictions are unreliable.
+We develop a geometric framework addressing both calibration and instance-level uncertainty quantification for neural network probability outputs.
+Treating probability vectors as points on the (c1)(c-1)-dimensional probability simplex equipped with the Fisher–Rao metric, we construct: (i) Additive Log-Ratio (ALR) calibration maps that reduce exactly to Platt scaling for binary problems while extending naturally to multi-class settings, and (ii) geometric reliability scores that translate calibrated probabilities into actionable uncertainty measures, enabling principled deferral of ambiguous predictions to human review.
+Theoretical contributions include: consistency of the calibration estimator at rate $O_{p}(n^{-1/2})$ via M-estimation theory (Theorem 1), and tight concentration bounds for reliability scores with explicit sub-Gaussian parameters enabling sample size calculations for validation set design (Theorem 2).
 We conjecture Neyman–Pearson optimality of our neutral zone construction based on connections to Bhattacharyya coefficients.
-Empirical validation on Adeno-Associated Virus classification demonstrates that the two-stage framework (calibration followed by reliability-based deferral) captures 72.5% of errors while deferring 34.5% of samples.
-Notably, this operational gain is achievable with any well-calibrated probability output; the contribution of geometric calibration lies in its theoretical foundations rather than empirical superiority over simpler alternatives.
-This work bridges information geometry and statistical learning, offering formal guarantees relevant to applications requiring rigorous validation.
+Empirical validation on Adeno-Associated Virus classification demonstrates that the two-stage framework captures 72.5% of errors while deferring 34.5% of samples, reducing automated decision error rates from 16.8% to 6.9%.
+Notably, calibration alone yields marginal accuracy gains; the operational benefit arises primarily from the reliability scoring mechanism, which applies to any well-calibrated probability output.
+This work bridges information geometry and statistical learning, offering formal guarantees for uncertainty-aware classification in applications requiring rigorous validation.
 Keywords: Calibration, uncertainty quantification, Fisher-Rao geometry, manifold geometry, probability simplex, geometric reliability scores, neural networks, neutral zones
-Preprint Note: This manuscript is under preparation for journal submission. Section 5 presents open problems where we invite community collaboration. Comments welcome at soumojit.das@wsu.edu.
+Preprint Note: This manuscript is under preparation for journal submission. Comments welcome at soumojit.das@wsu.edu.
 1 Introduction
@@
 This forced classification without uncertainty quantification leads to silent failures—the system appears to function normally while making unreliable predictions with potentially severe consequences.
-The fundamental issue is miscalibration: neural network probability outputs do not reflect true likelihood of correctness.
+The fundamental issue is twofold: neural network probability outputs are miscalibrated (they do not reflect true likelihood of correctness) and they lack instance-level uncertainty quantification (they provide no mechanism to identify which specific predictions are unreliable).
 Guo et al.
 [1] demonstrated that modern deep neural networks exhibit Expected Calibration Error (ECE) often exceeding 10% on standard benchmarks.
-This manifests as systematic overconfidence, where networks assign high probabilities to misclassified samples [2], and complete absence of uncertainty awareness [3].
+Yet even perfectly calibrated probabilities offer only population-level guarantees: knowing that 70% of predictions with confidence 0.7 are correct does not identify which 30% will fail.
+This gap between calibration and actionable uncertainty quantification leads to silent failures—systems appear confident while making unreliable predictions with potentially severe consequences [2, 3].
 1.1 Motivation: High-Stakes Applications
@@
 a Riemannian manifold equipped with the Fisher-Rao metric (Definition 1).
-This geometric perspective reveals that existing calibration methods—which apply transformations treating probability vectors as unconstrained Euclidean points—fundamentally misspecify the problem by ignoring the curved structure of probability space.
-Central theoretical result.
-Our main contribution is establishing that calibration via Additive Log-Ratio (ALR) transformations reduces exactly to Platt scaling for binary problems (Proposition 1).
-This positions geometric calibration as the natural multi-class generalization of a widely-used method, providing principled justification where none previously existed.
-The ALR framework respects the compositional structure of probability vectors [13] while enabling standard optimization techniques in the transformed space.
+This geometric perspective enables a unified treatment of both calibration and uncertainty quantification, addressing the two problems identified above through a coherent mathematical framework.
+Two-stage framework.
+We decompose uncertainty-aware classification into complementary stages:
+  1. 1.
+    Calibration (population-level): Affine transformation in ALR space corrects systematic probability distortions, reducing exactly to Platt scaling for binary problems (Proposition 1) while extending naturally to multi-class settings.
+  2. 2.
+    Reliability scoring (instance-level): Geometric reliability scores based on Fisher-Rao distance to simplex vertices translate calibrated probabilities into actionable uncertainty measures.
+Neutral zones defer predictions with low reliability to human review.
+This separation clarifies their distinct roles: calibration ensures probability estimates are accurate on average, while reliability scoring identifies which specific predictions warrant trust.
+The operational gains arise primarily from the second stage.
 Statistical guarantees.
@@
   2. 2.
-    Concentration (Theorem 2): Reliability scores based on Fisher-Rao distance satisfy tight concentration bounds with explicit sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$.
-These bounds enable sample size calculations for validation set design prior to data collection.
-Geometric neutral zones.
-Complementing calibration, we define reliability scores based on Fisher-Rao distance to the predicted simplex vertex (Definition 3) and construct neutral zones for principled deferral of uncertain predictions.
-The connection to Bhattacharyya coefficients (Remark 6) suggests potential Neyman-Pearson optimality, which we state as Conjecture 1.
-Importantly, the neutral zone mechanism is method-agnostic: it can be applied to any calibration method.
+    Concentration (Theorem 2): Reliability scores satisfy tight concentration bounds with explicit sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$.
+These bounds enable sample size calculations for validation set design prior to data collection—critical for applications where labeled validation data is expensive.
+Method-agnostic reliability scoring.
+The neutral zone mechanism can be applied to any calibration method.
 Our empirical analysis (Section 4.8, Table 5) confirms that operational gains are similar across calibration approaches.
-The contribution of geometric calibration is providing theoretical foundations, not empirical superiority.
+The contribution of geometric calibration is therefore twofold: (i) providing a principled multi-class generalization of Platt scaling grounded in information geometry, and (ii) enabling theoretical guarantees (Theorems 12) that simpler methods lack.
+The connection to Bhattacharyya coefficients (Remark 6) suggests potential Neyman-Pearson optimality (Conjecture 1).
 Empirical validation.
@@
   • Validation sample size ($n=310$) providing $\approx 5\times$ safety margin over theoretical requirement from Theorem 2
-The modest scale of this application is appropriate for proof of concept; validation on larger-scale and more diverse settings remains future work (Section 5).
+Notably, calibration alone improves accuracy only marginally (83.2% $\to$ 83.5%); the substantial error reduction comes from the reliability-based deferral mechanism.
+The modest scale of this application is appropriate for proof of concept; validation on larger-scale and more diverse settings remains future work.
 1.4 Novelty
 This work makes the following contributions:
-Mathematical framework:
+Instance-level uncertainty quantification:
+While calibration methods improve probability estimates on average, they do not identify which specific predictions are reliable.
+Our geometric reliability scores fill this gap, translating calibrated probabilities into actionable uncertainty measures.
+The framework enables principled deferral: predictions with low reliability are routed to human review, while high-reliability predictions proceed automatically.
+Empirically, this mechanism—not calibration itself—drives the operational gains (Table 4).
+Unified two-stage framework:
+We decompose uncertainty-aware classification into: (i) calibration (population-level probability correction, Section 2.2) and (ii) reliability scoring (instance-level uncertainty quantification, Section 4.5).
+This separation clarifies that calibration alone provides modest accuracy gains, while the combination with neutral zones (Section 4.6) achieves substantial error reduction.
+Our empirical analysis (Section 4.8, Table 5) demonstrates this is true regardless of calibration method, isolating the neutral zone mechanism’s contribution.
+Geometric foundations for multi-class calibration:
 To our knowledge, this is the first systematic application of Fisher-Rao geometry to post-hoc neural network calibration.
@@
 The key result—that geometric calibration reduces to Platt scaling for $c=2$ (Proposition 1)—validates our approach as a principled generalization rather than an arbitrary alternative.
-Theoretical guarantees:
+Theoretical guarantees with explicit constants:
 Theorem 2 provides the first concentration bounds for geometric reliability scores with explicit sub-Gaussian parameters.
 Previous work on uncertainty quantification (conformal prediction [9], deep ensembles [11]) either lacks sample complexity bounds or provides only asymptotic guarantees.
-Our explicit constants (Remark 10) enable practitioners to compute required validation set sizes.
-Unified two-stage framework:
-We decompose uncertainty-aware classification into: (i) calibration (population-level probability correction, Section 2.2) and (ii) reliability scoring (instance-level uncertainty quantification, Section 4.5).
-This separation clarifies that calibration alone provides modest accuracy gains (Table 4), while the combination with neutral zones (Section 4.6) achieves substantial error reduction.
-Our empirical analysis (Section 4.8, Table 5) demonstrates this is true regardless of calibration method, isolating the neutral zone mechanism’s contribution.
+Our explicit constants (Remark 10) enable practitioners to compute required validation set sizes before data collection.
 Information-theoretic grounding:
@@
 Section 4 provides empirical validation on AAV data.
-Section 5 discusses open problems and future directions.
+Section 5 discusses limitations, future directions, and broader context.
 2 Mathematical Framework
@@
 3 Theoretical Results
-We present three categories of results: proven theorems (Section 3.1), conjectures with strong evidence (Section 3.2), and research directions (Section 5).
+We present two categories of results: proven theorems (Section 3.1) and conjectures with strong evidence (Section 3.2).
 3.1 Proven Results
@@
 The theoretical rate $O_{p}(n^{-1/2})$ established in Theorem 1 remains the rigorous guarantee.
-Validation strategy for future work: To rigorously test whether super-efficiency occurs, we propose: (i) simulation studies with known $T^{*}$ where the true calibration map is specified, (ii) computing effective dimension via $d_{\text{eff}}=\operatorname{tr}(\nabla^{2}\mathcal{L}^{*})/\|\nabla^{2}\mathcal{L}^{*}\|_{2}$, and (iii) varying regularization strength $(\lambda_{1},\lambda_{2})$ to assess persistence of the phenomenon.
 3.2.2 Neyman-Pearson Optimality
@@
 […]
-Path to proof:
-  1. 1.
-    Binary case ($c=2$): For two classes, the Fisher-Rao distance reduces to $d_{\text{FR}}(\mathbf{p},\mathbf{e}_{1})=2\arccos(\sqrt{p_{1}})$, which is a monotone function of $p_{1}$.
-Under a Dirichlet noise model $\mathbf{p}^{\text{CNN}}\sim\text{Dir}(\alpha\cdot\mathbf{p}^{\text{true}})$ with concentration $\alpha>0$, the conditional density $f(d_{\text{FR}}\mid Y=1)$ versus $f(d_{\text{FR}}\mid Y=0)$ can be computed explicitly via change of variables from the Beta distribution.
-The likelihood ratio $f(d_{\text{FR}}\mid Y=1)/f(d_{\text{FR}}\mid Y=0)$ is then a function of $d_{\text{FR}}$ alone.
-We are currently deriving this density to verify monotone likelihood ratio analytically for $c=2$.
-  2. 2.
-[…]
-  3. 3.
-    General proof: The main obstacle is characterizing the distribution of $d_{\text{FR}}$ under Dirichlet perturbations on $\Delta^{c-1}$.
-Two potential approaches: (a) derive explicit formulas for geodesic distances under wrapped normal or Riemannian normal distributions on the simplex–these are known for certain symmetric cases [14], or (b) use concentration of measure: the positive curvature $K=+1/4$ implies Lévy-type concentration on spherical caps, which may yield sufficient control on conditional densities to establish MLR via comparison arguments.
-Spherical concentration is actually stronger than hyperbolic concentration for our purposes, as geodesic balls on positively curved spaces have smaller volume than Euclidean balls of the same radius.
-Validation strategy: Implement Monte Carlo validation for $c\in\{2,3,4,5\}$ with varying concentration parameters $\alpha\in\{1,5,10,50\}$.
-For each $(c,\alpha)$ pair, generate $10^{5}$ samples and estimate likelihood ratio $\widehat{LR}(d_{\text{FR}})=\hat{f}(d_{\text{FR}}\mid\text{correct})/\hat{f}(d_{\text{FR}}\mid\text{error})$ via kernel density estimation.
-Check monotonicity by testing whether $\partial\widehat{LR}/\partial d_{\text{FR}}<0$ holds across the support.
 4 Empirical Validation
@@
 However, for settings requiring formal validation documentation—regulatory submissions, safety-critical deployments, or expensive validation studies where sample size must be justified a priori—the theoretical foundations of geometric calibration provide value that empirical benchmarks alone cannot capture.
-5 Open Problems and Future Directions
-This section outlines research directions extending the geometric calibration framework.
-We organize by theme, progressing from theoretical extensions to practical questions and connections with related methods.
-5.1 Curvature-Based Error Bounds
-The positive curvature $K=+1/4$ of the probability simplex suggests tighter error propagation than Euclidean space.
-Specifically, geodesic triangles on positively curved spaces have angle sums exceeding $\pi$, and comparison theorems (Toponogov, Rauch) bound distance distortions.
-We conjecture:
-$d_{\text{FR}}(\mathbf{p},\mathbf{q})\leq 2\tanh^{-1}\left(\frac{\|\operatorname{ALR}(\mathbf{p})-\operatorname{ALR}(\mathbf{q})\|_{2}}{C}\right)$ (18)
-for some constant $C>0$, with the sublinear $\tanh^{-1}$ growth suggesting tighter bounds than linear Euclidean propagation.
-The intuition is that the ALR transformation “stretches” distances near the simplex boundary, while the spherical geometry “compresses” them, potentially yielding favorable cancellation.
-5.2 Distribution Shift Robustness
-Under covariate shift $Q\ll P$ with bounded likelihood ratio $dQ/dP\leq M$, we aim to prove:
-$\left|\operatorname{\mathbb{E}}_{Q}[d_{\text{FR}}]-\operatorname{\mathbb{E}}_{P}[d_{\text{FR}}]\right|\leq\sqrt{(M-1)^{2}+2M\cdot D_{KL}(Q\|P)}\cdot\pi$ (19)
-using the bounded diameter $\operatorname{diam}(\Delta^{c-1})=\pi$.
-This would provide guarantees for calibration transfer across related domains (e.g., training on one imaging modality, deploying on another).
-5.3 Geometry vs. Temperature Scaling: Phase Transitions
-Temperature scaling uses one parameter; geometric calibration uses $(c-1)^{2}+(c-1)$.
-We hypothesize a phase transition: geometric methods dominate when miscalibration is anisotropic, quantified by effective rank:
-$r_{\text{eff}}=\frac{\operatorname{tr}(\Sigma)}{\|\Sigma\|_{2}},\quad\Sigma=\operatorname{Cov}[\operatorname{ALR}(\mathbf{p}^{\text{CNN}})].$ (20)
-When $r_{\text{eff}}\approx 1$, miscalibration is nearly isotropic (all directions equally affected), and temperature scaling suffices.
-When $r_{\text{eff}}\gg 1$, different ALR directions require different corrections, favoring the full geometric approach.
-Characterizing the critical threshold $r^{*}_{\text{eff}}(c)$ as a function of number of classes remains open.
-5.4 Neutral Zone Size Reduction
-Does calibration reduce neutral zone size exponentially? Via Fano’s inequality, we conjecture:
-$|NZ^{\text{cal}}|/|NZ^{\text{uncal}}|\stackrel{{\scriptstyle?}}{{\approx}}\exp(-\theta\cdot\Delta\text{ECE})$ (21)
-for some $\theta>0$ depending on the class structure.
-The intuition is that improved calibration concentrates predictions away from decision boundaries, reducing the proportion requiring deferral.
-Empirical validation across diverse datasets would clarify the functional form.
-5.5 Extensions to r-Power Methodology
-The following directions extend the r-power framework [12] and will be developed in subsequent work.
-5.5.1 Multi-Class Extension
-Extension beyond binary classification requires testing ordered hypotheses via Union-Intersection principles. Specifically, we test:
-$H_{0}^{(j)}:p_{j}\leq\max_{k\neq j}p_{k}\quad\text{vs.}\quad H_{1}^{(j)}:p_{j}>\max_{k\neq j}p_{k}$ (22)
-for the predicted class $j=\hat{j}$.
-Our current approach uses step-down methods, adjusting for conservativeness through controlled false negative rates.
-The key challenge is maintaining power while controlling familywise error across the c(c1)/2c(c-1)/2 pairwise comparisons implicit in the multi-class setting.
-5.5.2 ADEBO Adaptation for Bounded Support
-The ADEBO estimator [24] requires approximate normality.
-For probability vectors with bounded support $\epsilon,1-\epsilon]^{c}\cap\Delta^{c-1}$, the logit transformation induces skewness near boundaries.
-Deriving moment conditions for bounded support—particularly higher-order corrections to the normal approximation—would extend ADEBO’s applicability to calibrated probability outputs.
 5 Conclusion
-We have developed a geometric framework for uncertainty-aware classification that bridges information geometry and statistical learning theory.
-The framework comprises two stages: geometric calibration of neural network probability outputs via ALR transformations, followed by reliability scoring and neutral zone construction for principled deferral of uncertain predictions.
+We have developed a geometric framework for uncertainty-aware classification that addresses both calibration and instance-level uncertainty quantification.
+The key insight is that calibration alone—even perfect calibration—provides only population-level guarantees, leaving unanswered the operationally critical question: which specific predictions should we trust?
+Our two-stage framework answers this by combining geometric calibration with reliability scoring that translates calibrated probabilities into actionable uncertainty measures.
 Main contributions.
-Our central theoretical result is Proposition 1, which establishes geometric calibration as the natural multi-class generalization of Platt scaling—a widely-used binary method that previously lacked principled extension to $c>2$ classes.
-This fills a gap in the calibration literature by grounding multi-class calibration in the intrinsic geometry of the probability simplex.
-Supporting this framework, Theorem 1 provides consistency guarantees at rate $O_{p}(n^{-1/2})$ via M-estimation theory, while Theorem 2 yields concentration bounds for reliability scores with explicit sub-Gaussian parameters.
-These bounds enable practitioners to determine required validation set sizes before data collection—a practical benefit for applications where labeled validation data is expensive.
+The primary operational contribution is the reliability scoring mechanism, which enables principled deferral of uncertain predictions.
+On AAV classification, this mechanism captures 72.5% of errors while deferring 34.5% of samples—performance that is largely independent of the specific calibration method used (Table 5).
+The geometric framework provides the theoretical foundations: Proposition 1 establishes geometric calibration as the natural multi-class generalization of Platt scaling, Theorem 1 provides consistency guarantees, and Theorem 2 yields concentration bounds with explicit sub-Gaussian parameters enabling sample size calculations before data collection.
 Empirical findings.
-Validation on AAV classification demonstrates that the two-stage framework achieves substantial error reduction: 72.5% error capture at 34.5% deferral rate, reducing automated decision errors from 16.8% to 6.9%.
-Importantly, our comparison with alternative calibration methods (Table 5) reveals that these operational gains arise primarily from the neutral zone mechanism itself, which is applicable to any well-calibrated probability output.
-Geometric calibration achieves competitive empirical performance while providing formal guarantees that simpler methods lack.
+Validation on AAV classification demonstrates that the two-stage framework achieves substantial error reduction: automated decision errors drop from 16.8% to 6.9%.
+Crucially, calibration alone improves accuracy only marginally (83.2% $\to$ 83.5%); the operational benefit arises from reliability-based deferral.
+This separation clarifies that while geometric calibration provides principled theoretical foundations, the neutral zone mechanism itself—applicable to any well-calibrated output—drives the practical gains.
 Limitations.
@@
 Future directions.
-The open problems in Section 5 outline extensions of this work.
-Most pressing are: (i) rigorous investigation of Conjecture 1, potentially via explicit characterization of the likelihood ratio under Dirichlet perturbations; (ii) distribution shift robustness bounds for calibration transfer across related domains; and (iii) integration with the r-power methodology [12] for distribution-free neutral zones in multi-class settings.
+Several open problems merit investigation: (i) rigorous proof of Conjecture 1 on Neyman-Pearson optimality; (ii) extension to high-dimensional settings ($c\gg 3$) and analysis under distribution shift; and (iii) integration with the r-power methodology [12] for distribution-free neutral zones in multi-class settings.
 The connection between Fisher-Rao geometry and optimal decision theory suggested by Remark 6 merits deeper exploration.
 Broader context.
-As AI systems increasingly inform high-stakes decisions in medicine, criminal justice, and autonomous systems, principled uncertainty quantification becomes essential.
-This work contributes formal foundations for one approach to this challenge.
-The theoretical guarantees established here—consistency rates, concentration bounds, and the principled connection to Platt scaling—may prove particularly valuable in applications requiring rigorous validation documentation, where empirical performance benchmarks alone are insufficient.
+As AI systems increasingly inform high-stakes decisions in medicine, criminal justice, and autonomous systems, the gap between calibration and actionable uncertainty quantification becomes critical.
+Well-calibrated probabilities tell us that 70% of predictions at confidence 0.7 are correct; they do not tell us which 30% will fail.
+This work contributes formal foundations for bridging that gap, offering instance-level reliability scores with theoretical guarantees that may prove particularly valuable in applications requiring rigorous validation documentation.
 Appendix A Proofs
@@
 Verification of Glivenko-Cantelli conditions.
+We start with verifying the conditions used in example 19.8 of van der Vaart.
   1. 1.